### v1

# Semantics-Preserving Evasion of LLM Vulnerability Detectors

Abstract
LLM-based vulnerability detectors are increasingly deployed in security-critical code review, yet their resilience to evasion under behavior-preserving edits remains poorly understood. We evaluate detection-time integrity under a semantics-preserving threat model by instantiating diverse behavior-preserving code transformations on a unified C/C++ benchmark (N=5000N{=}5000), and introduce a metric of joint robustness across different attack methods/carriers. Across models, we observe a systemic failure of semantic invariant adversarial transformations: even state-of-the-art vulnerability detectors perform well on clean inputs while predictions flip under behavior-equivalent edits. Universal adversarial strings optimized on a single surrogate model remain effective when transferred to black-box APIs, and gradient access can further amplify evasion success. These results show that even high-performing detectors are vulnerable to low-cost, semantics-preserving evasion. Our carrier-based metrics provide practical diagnostics for evaluating LLM-based code detectors.
Machine Learning, ICML
Figure 1: Why carrier-constrained optimization matters. Left: Carrier-constrained GCG via identifier substitution preserves compilation and flips the detector to BENIGN. Middle: Naively appending a GCG-optimized string breaks compilation, so it is not a valid semantics-preserving evasion. Right: Unoptimized semantics-preserving edits remain compilable but typically fail to evade detection. Red denotes the GCG-optimized identifier, blue denotes a random identifier, and boldface highlights the original identifier.
1 Introduction
Software vulnerabilities remain among the most exploited attack vectors, driving rapid adoption of automated detection in modern continuous integration pipelines (Verizon, 2025; Microsoft, 2025). LLM-based detectors are now integrated into code review workflows at scale. This raises a security-critical question: do such detectors satisfy semantic invariance, or specifically, can an attacker evade detection using perturbations where the transformed program remains strictly compilable and retains the underlying vulnerability? If the latter holds, vulnerable code could reliably slip through automated security checks via innocuous transformations, including renaming identifiers, inserting comments, injecting dead-branch code, or adding preprocessor-guarded code, while leaving the underlying exploit fully intact. Existing research has highlighted the broad spectrum of adversarial vulnerabilities in LLMs (Fu et al., 2024; Vassilev et al., 2025). Prior adversarial methods often ignore syntactic constraints, producing perturbations that break compilation or alter program logic (Ebrahimi et al., 2018). These invalid edits are practically useless: they cannot preserve the vulnerability’s executable behavior. As illustrated in Figure 1, effective evasion presents a dual challenge: naive GCG attack (Middle) breaks compilation, while randomized semantic edits (Right) often fail to evade detection. A valid attack must simultaneously satisfy the compiler’s strict syntax rules and the adversary’s evasion objective (Left).
More general-purpose LLM safety evaluations, on the other hand, primarily prioritize preventing harmful outputs, such as toxic content and jailbreak-style policy violations (Derczynski et al., 2024; Wei et al., 2025). Vulnerability detection, however, confronts an orthogonal integrity threat: adversaries need not induce malicious generation; they only need to make the detector oblivious to an existing vulnerability through behavior-preserving edits. This exposes a gap in current evaluation practices: standard safety probes emphasize output harmfulness, while clean-accuracy benchmarks ignore robustness under semantic-preserving transformations. Consequently, a detector can appear safe under toxicity-focused evaluations and achieve strong benchmark scores, yet remain brittle to innocuous perturbations in real-world deployment.
To bridge this gap, we present a systematic evaluation framework for detection-time integrity under a semantics-preserving threat model. We instantiate multiple behavior-equivalent attack surfaces and evaluate both open-weight and proprietary models (via surrogate transfer) on a unified C/C++ benchmark. Beyond per-attack evasion, we introduce complete resistance, measuring the fraction of vulnerabilities that withstand all semantics-preserving transformations. We make the following contributions:
  1. 1.
    First Gradient-Based Evaluation of Semantic Invariance under Carrier Constraints. We combine gradient-driven optimization with carrier-constrained perturbations to expose LLM detector vulnerabilities. Preprocessor-style carriers (#ifdef, macros) achieve high success rates via universal triggers, contrasting with search-based methods (Jiang et al., 2026). We introduce Complete Resistance (CR) to measure joint robustness across carriers.
  2. 2.
    Exposing the Fragility of High-Performing Detectors. We show that models achieving 70%+ clean accuracy collapse under minimal perturbations: fewer than 13% of vulnerabilities resist all carrier types (CR << 13%), meaning 87% of ”detected” threats can be evaded by at least one semantics-preserving edit (e.g., #ifdef insertion). High benchmark performance masks critical robustness failures. Detectors succeed on benign inputs but fail severely when code undergoes innocuous transformations.
  3. 3.
    High Exploitability of Black-box APIs via Low-barrier Transfer. We reveal that universal adversarial strings optimized on small surrogate models transfer effectively to closed-source models like GPT-4o without additional optimization.
2 Related Work
Datasets for Vulnerability Detection. Benchmarks for evaluating vulnerability detectors consist of labeled C/C++ functions sourced from real-world projects. Early datasets like BigVul and DiverseVul contain label noise and train-test leakage (Chen et al., 2023; Fan et al., 2020), while PrimeVul addresses these issues through manual curation and stricter deduplication (Ding et al., 2024). Since our adversarial evaluation requires clean labels-a mislabeled ”vulnerable” sample would falsely appear robust-we adopt PrimeVul’s protocols and extend them with carrier-specific controls (§4.1).
LLMs for Vulnerability Detection. Recent benchmarks, such as (Khare et al., 2025), report that state-of-the-art LLMs can achieve promising performance on standard datasets. However, other studies indicate that code language models remain unreliable for detection in realistic settings, paralleling similar fragility observed in code reasoning tasks (Orvalho & Kwiatkowska, 2025), highlighting the need to evaluate robustness beyond clean accuracy (Ding et al., 2024). While separate efforts benchmark the security of generated code (e.g., SafeGen-style evaluations), our focus is strictly on detection-time robustness under semantics-preserving edits, not on generation safety (Li et al., 2025).
Concurrently, recent work studies input-side defenses for adversarial transformations on code LMs. Evaluate-and-Purify proposes EP-Shield, using an LLM-as-a-judge to assess naturalness and purify identifier-substitution adversarial examples to restore predictions (Mu et al., 2025).
Our work is complementary: we focus on attack-centric robustness evaluation under a semantics-preserving threat model spanning identifier substitution, comment-carrier insertion, preprocessor-based insertions, and dead-branch code carrier, and report metrics (ASR, Complete Resistance, Transferability) that can also be used to quantify the gains of such purification wrappers across attack surfaces.
Code-Specific Adversarial Attacks on Vulnerability Detection. Adversarial attacks on code models span multiple representations. NatGVD (Rath et al., 2025) modifies graph structures for GNN detectors, while neural binary analyzers face syntactic mutation risks (Bundt et al., 2023). We introduce gradient-driven, carrier-constrained analysis to address this gap. Our universal triggers reveal that preprocessor carriers achieve 18%–99% ASR variance across models, fundamentally differing from identifier-based attacks. We propose Complete Resistance (CR) to measure joint robustness, a dimension orthogonal to SVulAttack’s instance-specific optimization.
Jailbreaks and Transferable Prompt Attacks. Adversarial attacks on aligned LLMs have evolved from manual template engineering to automated frameworks. Optimization-based methods, such as GCG (Zou et al., 2023) and AutoDAN (Zhu et al., 2023), utilize gradient-guided or genetic search to find transferable adversarial suffixes. Concurrently, LLM-guided black-box attacks like PAIR (Chao et al., 2024), decomposition attack (Brown et al., 2025), and TAP (Mehrotra et al., 2024) leverage attacker models to iteratively refine prompts. Recent studies have extended this by characterizing the “utility” of such outputs (Nikolić et al., 2025). Existing evaluations, however, rarely examine how such transferable prompts interact with code tasks in a detection setting where edits must preserve compilability and semantics. We adapt GCG to this constrained setting.
3 Threat Model & Attacks
3.1 Threat Model
Adversary objective. We study evasion attacks against LLM-based vulnerability detectors. Given a ground-truth VULNERABLE function $x$ and a detector $f(\cdot)$, the adversary aims to produce a perturbed function $x^{\prime}$ such that $f(x^{\prime})=\texttt{BENIGN}$ (a false negative), allowing the vulnerability to bypass automated security gates.
Capabilities and semantic-invariance constraint. The adversary may modify only the source-code input while aiming to preserve semantic invariance; the transformed program must remain compilable and retain the underlying vulnerability. We restrict edits to four carriers that are intended to be semantics-preserving by construction: (i) scope-consistent identifier substitution, (ii) non-executable comment carriers, (iii) preprocessor carriers (macros/conditional guards) that are inactive under our compilation setting (e.g., unused macro definitions or disabled guards), and (iv) dead-branch insertion with a statically false predicate. We automatically validate carrier well-formedness (scope-consistent replacement, balanced comment delimiters, well-nested preprocessor directives/macro definitions, and syntactically valid dead-branch templates); we report pass rates and retain all samples without post-hoc filtering, validated by two-tier analysis: differential syntactic analysis (Tree-sitter (Khare et al., 2025)) on 5,000 functions and compilation-based validation (gcc) on another 1,434 standalone-compilable functions (Appendix A.3, Tables 67).
Knowledge regimes. We evaluate robustness under three realistic access assumptions:
  • White-box. The adversary has full access to model weights and gradients, enabling direct gradient-guided optimization (e.g., GCG) of adversarial strings.
  • Black-box (API) via transfer. Targeting proprietary models with inaccessible internals, the adversary optimizes universal strings on a smaller surrogate model and applies them to the target without further modification.
  • Operational black-box. Targeting large locally-hosted models where weights are accessible but direct gradient-based optimization is computationally expensive for typical adversaries. We primarily evaluate this regime using transfer-only methods to establish a realistic baseline of low-cost exploitability. However, to quantify the gap between this low-cost approach and the worst-case security risk, we explicitly perform an on-target white-box ablation (see Table 5) as a high-cost upper bound.
Detailed model specifications and versions are provided in Appendix A.1.
3.2 GCG-based Optimization
Code presents a unique constraint: syntax errors break compilation. Standard GCG generates random-looking strings that often violate language grammar. We address this by restricting optimization to four carrier types with automated validity checks.
We use the Greedy Coordinate Gradient (GCG) algorithm to optimize a discrete, printable adversarial string $\sigma$ that induces a prediction flip from VULNERABLE to BENIGN. GCG uses token-level gradients to guide coordinate-wise updates in a discrete search space, while all edits are restricted to the semantics-preserving transformation space defined by our threat model. We choose GCG for its ability to produce a universal string that transfers across both samples and models (§5.4).
Optimization Objective and Universal Strategy. The attacker maximizes the model’s relative preference for the target label BENIGN over the ground truth VULNERABLE. Specifically, for a code instance $x$ and carrier $\mathcal{T}(x,\sigma)$, we maximize the log-odds objective $\mathcal{J}(\sigma;x)=\log p_{\theta}(\texttt{BENIGN}\mid\mathcal{T}(x,\sigma))-\log p_{\theta}(\texttt{VULNERABLE}\mid\mathcal{T}(x,\sigma))$. To ensure generalization, rather than optimizing per-instance, we learn a single universal adversarial string $\sigma$ over a fixed support set $\mathcal{X}_{\text{opt}}$ (m=10m{=}10) by maximizing the mean performance: $\mathcal{J}_{\text{univ}}(\sigma)=\frac{1}{m}\sum_{x\in\mathcal{X}_{\text{opt}}}\mathcal{J}(\sigma;x)$. Post-optimization, $\sigma$ is frozen to evaluate zero-shot transferability. The success of such input-agnostic triggers suggests that the detector’s vulnerability stems from systematic feature bias rather than instance-specific brittle directions.
Implementation details. Optimization hyperparameters, surrogate model selection (Qwen2.5-Coder-14B), and prompt templates are provided in Appendix B.5 and Appendix E.
3.3 Attack Categories
We define the transformation operator $\mathcal{T}(x,\sigma)$ through four families of insertion carriers. The following methods ensure functional equivalence by construction, modifying only the source-code input while strictly maintaining compilability and executable behavior (see Appendix F for details):
  1. 1.
    Identifier substitution. We apply a consistent renaming map $\mathcal{M}:v\mapsto\sigma$ to the specific identifier $v$ explicitly flagged by the detector (i.e., extracted from the "variable" JSON field), or to the most frequent local identifier if attribution is unavailable. We replace all occurrences of $v$ within its scope. This focuses the perturbation on lexical cues without affecting variable binding or logic (avoiding variable shadowing).
  2. 2.
    Comment-carrier insertion. We inject $\sigma$ into syntactically valid comment carriers (e.g., prepending a multi-line header). Comments act as transparent containers for adversarial tokens, allowing the insertion of non-interpretable strings.
  3. 3.
    Preprocessor-based insertion (macro/guard carriers). We inject $\sigma$ into inactive preprocessor regions, specifically unused macro definitions (denoted macro_name) or disabled conditional guards. These regions are stripped or ignored during compilation, ensuring zero impact on runtime behavior.
  4. 4.
    Dead-branch code insertion. We inject $\sigma$ into a formally unreachable control-flow block (e.g., if (0) { }). Unlike preprocessor directives, this carrier is syntactically indistinguishable from executable code, serving as a diagnostic for the detector’s ability to resolve basic Boolean control flow.
Family-specific candidate constraints. To ensure GCG-optimized strings remain valid under each carrier, we apply family-specific constraints to the candidate alphabet. For identifier substitution and dead-branch code, we enforce standard C-identifier compliance (alphanumeric with underscores, non-numeric start). For macro bodies and preprocessor-guarded blocks, we restrict the search space to a printable subset and enforce single-line constraints to prevent preprocessor syntax errors. Detailed vocabulary definitions and filtering rules are provided in Appendix A.2.
Default carriers and ablations. Unless otherwise specified, comment-carrier insertion uses a leading/header carrier, and macro insertion uses the macro-name carrier (macro_name). We study alternative placements (middle/trailing) and the macro-body carrier (macro_body) only in a focused ablation (Table 9).
4 Experimental Setup
4.1 Datasets Construction
Unified benchmark and design goals. To isolate robustness to semantics-preserving edits from confounders such as duplication leakage and length artifacts, we construct a single unified benchmark, Unified-VUL-N (N=5000N{=}5000). We pool vulnerable functions from PrimeVul, BigVul, and DiverseVul, enforcing three strict controls: (i) global cross-source de-duplication to prevent inflation from repeated functions, (ii) a fixed source composition to standardize evaluation across models, and (iii) a shared prompt budget to ensure valid comparisons.
Deduplication by normalization-only hashing. To prevent evaluation inflation from surface-level duplicates, we canonicalize code logic (stripping comments and formatting) to define a deterministic equivalence class via SHA-256 hashing. This normalization is used strictly for cross-source auditing; models are always evaluated on the original, unmodified source code. Detailed normalization rules are provided in Appendix B.1.
Filtering and Sampling. We restrict code to $\leq 4096$ tokens to reserve headroom for adversarial carriers and eliminate truncation-driven artifacts. A deterministic, quota-controlled sampling strategy ensures a stable and auditable benchmark composition (also see Appendix B.1).
4.2 Evaluation Protocol
Strict Parsing and Conservative Success Criterion. We employ a deterministic strict parser to ensure robustness measurements reflect true detection-time integrity rather than output formatting artifacts. A successful evasion is counted iff the model produces a valid JSON with label="BENIGN" for a ground-truth vulnerable input. Parsing failures, refusals, or hallucinations are counted as Resist (failure to evade). This yields a conservative estimate of attack success: the adversary wins only by forcing a decisive false negative. Empirically, output stability is high ($99.9\%$ valid parse rate), confirming that measured failures stem from semantic understanding gaps, not syntax degradation.
Evaluation Decomposition: Conditional Evasion. To isolate robustness from clean-accuracy capabilities, we evaluate attacks conditionally on Clean True Positives ($\mathrm{TP}_{\mathrm{clean}}$). Since a pre-existing error cannot be ”further evaded,” applying attacks to the entire dataset would inflate perceived robustness. We define $\mathrm{Flip}(A)$ as the subset of $\mathrm{TP}_{\mathrm{clean}}$ successfully evaded under attack $A$, and $\mathrm{Resist}(A)$ as those that maintain detection.
Primary Robustness Metrics. We report three metrics to characterize the Robustness & Accuracy Gap:
  1. 1.
    Conditional Attack Success Rate ($\mathrm{ASR}_{\mathrm{cond}}$). The fraction of correctly detected vulnerabilities that are flipped by a specific attack vector ($|\mathrm{Flip}|/|\mathrm{TP}_{\mathrm{clean}}|$).
  2. 2.
    Complete Resistance (CR). The fraction of vulnerabilities that withstand all evaluated attack families ($\cap_{k}\mathrm{Resist}(A_{k})/|\mathrm{TP}_{\mathrm{clean}}|$). This serves as the security floor, indicating the proportion of reliable detections that are truly invariant to semantic perturbations.
  3. 3.
    End-to-End Recall Drop ($\Delta\mathrm{TPR}$). The absolute decrease in detection coverage on the full dataset ($|\mathrm{Flip}_{\cup}|/N$). This measures the operational risk of deploying the model in an adversarial environment.
Formal derivations and single-family variants are detailed in Appendix B.2.
Baseline Clean Coverage. All attack metrics are contextualized against the clean performance baselines reported in Table 1, establishing the performance ceiling from which robustness degrades.
Table 1: Baseline clean coverage on the unified C/C++ evaluation set. Absolute counts and $\mathrm{TPR}_{\mathrm{clean}}$ are reported; attack metrics ($|\mathbf{Flip}_{\mathrm{union}}|$, $|\mathbf{Resist}_{\mathrm{complete}}|$) appear in §5.1.
Model NN $|\mathbf{TP}_{\mathrm{clean}}|$ $|\mathbf{FN}_{\mathrm{clean}}|$ $\mathrm{TPR}_{\mathrm{clean}}$ (%)
Qwen2.5-Coder-32B 5000 1121 3879 22.42
Llama3.1-8B 5000 1552 3448 31.04
CodeAstra 5000 3454 1546 69.08
GPT-4o 5000 2300 2700 46.00
GPT-5-mini 5000 3681 1319 73.62
5 Results
5.1 Main Results: Overall Attack Effectiveness
One would expect that a detector achieving 70% clean accuracy provides reliable security coverage for the vulnerabilities it identifies. Our evaluation reveals the opposite: across all models except GPT-5-mini, over 87% of correctly detected vulnerabilities can be flipped to false negatives using only semantics-preserving edits (Figure 2). CodeAstra, despite correctly flagging 3,454 vulnerabilities on clean inputs, maintains robust detection for only 4 functions under on-target attacks (CR = 0.12%), a 99.88% collapse in detection integrity.
This pattern occurs under realistic access assumptions: surrogate-transfer attacks on black-box APIs (GPT-4o, GPT-5-mini) and operational black-box models (Qwen2.5-Coder-32B), supplemented by on-target optimization for accessible models (Llama3.1-8B, CodeAstra) to establish upper bounds. Across this threat landscape, GPT-5-mini is an exception, exhibiting substantially lower attack success (44.04% union $\mathrm{ASR}_{\mathrm{cond}}$) compared to other models (87%–100%), suggesting stronger invariance to carrier-based perturbations. These results indicate that clean benchmark accuracy is not a security guarantee. An attacker needs only to evade the subset of cases a detector would otherwise catch, and our results show this is achievable through innocuous transformations like identifier renaming or comment insertion (Table 2; clean baseline in Table 1).
Figure 2: Overall attack effectiveness across models. Solid bars: transfer attacks using universal strings from Qwen2.5-Coder-14B surrogate. Hatched bars: on-target optimization (Llama3.1-8B, CodeAstra only). (a) Union $\mathrm{ASR}_{\mathrm{cond}}$ across carriers. (b) End-to-end recall decrease ($\Delta\mathrm{TPR}$).
Benign-class invariance. To verify that our edits induce targeted false negatives rather than indiscriminate label corruption, we evaluate attacks on a held-out set of 500 ground-truth benign functions. Across all models and attack methods, we observe 100% benign-class invariance: every function predicted as BENIGN under clean inference remains BENIGN after applying our transformations. This supports that the observed flips reflect targeted integrity violations rather than indiscriminate label corruption.
5.2 Carrier-Specific Vulnerabilities: Preprocessor Attacks Dominate
Table 2: Conditional attack success rate ($\mathrm{ASR}_{\mathrm{cond}}$) under surrogate transfer. Universal strings from Qwen2.5-Coder-14B applied to all targets.
Model Identifier Sub. Comment ifdef macro dead-branch
GPT-4o 31.09% 65.48% 71.61% 42.35% 19.09%
GPT-5-mini 10.62% 18.34% 28.14% 21.57% 17.58%
Qwen2.5-Coder-32B 26.94% 91.53% 98.84% 85.81% 63.15%
Llama3.1-8B 34.66% 64.50% 99.03% 96.33% 59.79%
CodeAstra 4.72% 78.60% 42.59% 75.48% 72.12%
We break down attack effectiveness by carrier under surrogate transfer (Table 2). Three patterns emerge. First, no single carrier dominates: ifdef achieves near-complete evasion on Llama3.1-8B (99.03%) and Qwen2.5-Coder-32B (98.84%) but achieves moderate success elsewhere. Second, preprocessor carriers often outperform identifier substitution, suggesting over-reliance on non-executable boundary tokens. Third, GPT-5-mini exhibits uniformly low ASR (10–28%), indicating broader invariance. Besides, Dead-branch carriers are also effective on several models, indicating limited dead code alignment.
Model-specific patterns. Most models show high vulnerability to preprocessor carriers, with Llama3.1-8B and Qwen2.5-Coder-32B near-ceiling (85–99%). CodeAstra presents a distinct profile: identifier substitution transfers poorly (4.72%), while boundary carriers remain effective (42–79%). On-target optimization reverses this pattern (§5.4).
Positional and functional sensitivity. Attack success is sensitive to carrier placement and syntactic role. Trailing comment carriers outperform leading ones for Llama3.1-8B and CodeAstra, consistent with boundary-position bias in long-context aggregation. Similarly, macro-name carriers are considerably more potent than macro-body carriers (detailed ablation in Table 9). Function length shows weak correlation with evasion (Pearson $r\in[-0.17,-0.02]$), and CWE-level analysis reveals modest variation (Appendix D.1 and Figure 4).
5.3 Sanitization Eliminates Carriers but Induces Prediction Drift
Table 3: Sanitization eliminates targeted carriers (ASR $\rightarrow$ 0%) but induces prediction drift on clean inputs (see Table 10 for details).
[…]
Qwen2.5-Coder-32B 91.53→0% 98.84→0% 85.81→0% +5.2%
Llama3.1-8B 64.50→0% 99.03→0% 96.33→0% +8.6%
CodeAstra 78.60→0% 42.59→0% 75.48→0% +0.1%
Given the high success of comment- and preprocessor-based carriers, we evaluate a simple input-sanitization baseline that removes these carrier surfaces prior to inference. The pipeline (i) strips C/C++ line and block comments, and (ii) canonicalizes/removes preprocessor directives before passing code to the detector. We treat this as a carrier-elimination baseline rather than a semantics-aware defense.
Table 3 shows that sanitization successfully drives ASR to zero for targeted carriers. However, it induces substantial prediction drift: stripping benign comments and preprocessor directives from unperturbed, ground-truth vulnerable functions increases clean detection rates by up to 8.6% (Llama3.1-8B: 1552→1685 true positives). This reveals that models utilize non-executable lexical context as decision-relevant features, indicating a lack of semantic invariance even absent adversarial perturbations (detailed drift analysis in Table 10, Appendix D).
Moreover, sanitization is not a free lunch: comments and preprocessor directives appear in 57.0% and 12.8% of functions, respectively, meaning this defense rewrites a substantial fraction of inputs. We exclude dead-code elimination, as it risks mis-eliminating live code without compiler-level context (Tu et al., 2024). Furthermore, sanitization fails to address executable-looking carriers like identifier substitution and dead-branch insertions.
5.4 Transfer Attacks Succeed on Black-Box APIs; On-Target Reveals Complex Interactions
Table 2 presents transfer results where universal strings optimized on Qwen2.5-Coder-14B are applied unchanged to all targets. This models a low-barrier adversary: optimization is performed once, and resulting strings are reused across vulnerable functions and deployed detectors. Across black-box and operational black-box targets, transfer achieves 44.04%–99.03% union $\mathrm{ASR}_{\mathrm{cond}}$, establishing that API opacity alone does not prevent evasion under semantics-preserving perturbations. However, transferability is not uniform: GPT-5-mini exhibits materially lower success (44.04%) than other models (87–99%), indicating that robustness varies across model families even under identical carriers.
On-target optimization reveals divergent vulnerability profiles. While transfer attacks establish a practical lower bound on exploitability, they may underestimate the true vulnerability of a model. To quantify the gap between low-cost transfer and worst-case security risk, we perform on-target universal GCG for Llama3.1-8B and CodeAstra (accessible models). Table 4 compares transfer and on-target regimes across carriers.
Table 4: Transfer vs. on-target attack success. Transfer uses universal strings from Qwen-14B surrogate; on-target optimizes directly on the target.
Model Regime $\mathrm{ASR}_{\mathrm{cond}}$ (%)
Idsub Comment ifdef macro Dead-br
Llama3.1-8B Transfer 34.66 64.50 99.03 96.33 59.79
On-target 43.49 42.01 99.03 65.46 86.66
CodeAstra Transfer 4.72 78.60 42.59 75.48 72.12
On-target 85.41 63.06 16.62 98.87 97.51
For Llama3.1-8B, on-target optimization improves identifier substitution (34.66% $\rightarrow$ 43.49%) and dead-branch insertion (59.79% $\rightarrow$ 86.66%), but degrades comment (64.50% $\rightarrow$ 42.01%) and macro (96.33% $\rightarrow$ 65.46%). This suggests that universal strings optimized on a surrogate can outperform target-specific optimization for certain carriers, likely because surrogate-optimized universal strings exploit more transferable boundary-token sensitivities.
CodeAstra exhibits a clear asymmetry. Transfer achieves 4.72% identifier success but 42.59% for ifdef. On-target optimization inverts this: identifier reaches (85.41%) while ifdef collapses (16.62%).
We hypothesize two mechanisms. For identifier substitution, CodeAstra’s loss landscape provides smooth gradients: GCG efficiently discovers variable-name perturbations that flip predictions. For ifdef, on-target GCG likely encounters optimization difficulty, so the optimizer converges to poor local minima that fail to exploit the vulnerabilities that surrogate-optimized strings (accidentally) discovered through multi-model training. This suggests that gradient-based optimization success depends not only on model vulnerabilities but also on the optimizability of the attack surface: some carrier spaces may be inherently harder to optimize even when vulnerabilities exist.
Qwen2.5-Coder-32B: Surrogate capacity and on-target upper bound. For Qwen2.5-Coder-32B (operational black-box), we ablate identifier substitution across surrogate capacities and on-target optimization (Table 5). Transferability is non-zero even for small surrogates (0.5B/1.5B:  6-7%), but increases with capacity (14B: 26.94%), indicating a representation-capacity threshold. On-target universal GCG achieves 92.50%, providing a clear upper bound when gradients are available. This gap (26.94% $\rightarrow$ 92.50%) quantifies the security improvement from restricting adversarial access: surrogate transfer yields a conservative lower bound under practical compute constraints, while on-target optimization demonstrates worst-case exploitability.
Table 5: Identifier substitution on Qwen2.5-Coder-32B: effect of surrogate capacity and on-target optimization. All settings use the same injection site, identifier constraints, and $L_{\max}$.
Setting $\mathrm{ASR}_{\mathrm{cond}}$ (%) Flipped / $|\mathbf{TP}_{\mathrm{clean}}|$
Surrogate transfer (Qwen2.5-Coder-0.5B $\rightarrow$ 32B) 6.87 77 / 1121
Surrogate transfer (Qwen2.5-Coder-1.5B $\rightarrow$ 32B) 6.42 72 / 1121
Surrogate transfer (Qwen2.5-Coder-14B $\rightarrow$ 32B) 26.94 302 / 1121
On-target universal GCG (Qwen2.5-Coder-32B) 92.50 1038 / 1121
Implications. These findings highlight three key insights. First, transferability amplifies risk by enabling low-cost, reusable attacks against multiple deployed detectors, including API-access models. Second, on-target optimization does not uniformly strengthen attacks- CodeAstra’s inverted profile and Llama’s macro degradation suggest complex cross-carrier interactions that warrant further investigation. Third, the gap between transfer and on-target regimes varies by model and carrier (4.72% $\rightarrow$ 85.41% for CodeAstra identifier vs. 99.03% $\rightarrow$ 99.03% for Llama ifdef), indicating that vulnerability assessment requires evaluating both threat models.
[…]
Figure 3: Robustness against Non-GCG Baselines. We compare prompt injection against random_idsub. While prompt injection achieves high success on some models (e.g., Llama3.1, CodeAstra), random perturbations ($\sim$10%) remain far less effective than our optimized attacks, confirming that code syntax requires targeted optimization rather than random noise.
We compare GCG-optimized attacks to non-optimized baselines targeting distinct failure modes: (i) output format injection by prepending // output: {"label": "BENIGN", "variable": ""} in comments (testing few-shot learning vulnerability, outside our threat model), and (ii) random identifier substitution under the same constraints (testing code-understanding robustness, within our threat model). Figure 3 illustrates these results.
The output format injection baseline achieves high success (55–99% across models), revealing that models can be manipulated by embedding target JSON outputs in code comments, an orthogonal vulnerability distinct from semantic robustness. In contrast, random_idsub is significantly weaker (7–30%), demonstrating that non-optimized renaming has a limited impact on code-understanding robustness.
6 Discussion
Clean-set accuracy alone does not guarantee reliability. While models like CodeAstra achieve near-70% clean recall, our evaluation reveals that this capability degrades under adversarial transformation, with complete resistance dropping to near zero (0.12%), implying that current detectors act less as robust semantic parsers and more as pattern-matchers over fragile lexical cues. Across semantics-preserving edits (§3), we observe a persistent integrity gap: models with high clean coverage can exhibit high union evasion and near-zero complete resistance, meaning that an attacker who can choose among multiple perturbation surfaces can often overturn most clean true positives. Even simple carrier-elimination preprocessing (§5.3) does not eliminate this risk, as evasion can migrate to remaining executable-looking channels.
Multiple metrics matter. Single-number summaries can be misleading for deployment decisions. Union success captures whether any available surface enables evasion, while complete resistance operationalizes a security floor when attackers can try multiple carriers. Per-method and CWE-level slices then identify where robustness breaks (which channels, which vulnerability types), enabling targeted mitigation rather than blanket conclusions from aggregate ASR.
Temperature and transfer analysis. The observed fragility is structural rather than stochastic. Evasion rates remain stable across decoding temperatures (Appendix D.2), ruling out random sampling artifacts. More importantly, the high transferability of attacks from a 14B surrogate model to GPT-4o points to shared representational blind spots: distinct models appear to over-index on identical non-executable tokens (e.g., #ifdef guards) while ignoring the invariant executable logic.
Decoupling of Domain Capability and Robustness. The performance profile of CodeAstra reveals a meaningful insight: although task-specific fine-tuning boosts baseline detection, it does not guarantee robustness. CodeAstra exhibits high vulnerability in specific channels (e.g., macro insertion) despite high clean accuracy. This suggests that domain specialization can coexist with-or mask-structural fragility, emphasizing that fine-tuned models require integrity evaluations just as rigorous as general-purpose ones.
Implications for Deployment and Defense. Our findings suggest three concrete directions for hardening detectors. First, training on carrier-diverse adversarial examples (spanning identifier, comment, and preprocessor surfaces) may enforce broader invariance than single-method defenses. Second, Complete Resistance should join clean accuracy as a co-equal deployment metric- a detector is only as robust as its weakest carrier. Third, ensemble approaches combining sanitization (for boundary carriers) with robust base models may offer defense-in-depth where neither alone suffices.
Limitations and Future Directions. We focus on C/C++ functions with a fixed prompting protocol; extending to other languages and detection formulations would strengthen generality. While Tier 2 validation (differential AST analysis) scales to 5,000 functions and Tier 3 (compilation-based) provides strong proof on 1,434 functions (Appendix A.3), absolute behavioral equivalence requires full execution environments unavailable in function-level datasets. Future work should explore adversarial fine-tuning on semantics-preserving carriers, though computational cost and overfitting risks motivate investigating lightweight defense wrappers (Madry et al., 2019; Varshney et al., 2024).
7 Conclusion
On a unified C/C++ benchmark (N=5000N{=}5000), we present an integrity evaluation of LLM-based vulnerability detectors under semantics-preserving code perturbations, evaluating five carriers spanning four insertion families (identifier substitution, comments, preprocessor insertions, and dead-branch insertion) and quantifying robustness via union success and complete resistance. Across models, we observe widespread failures of semantic invariance: many vulnerabilities that are correctly detected on clean inputs can be flipped to false negatives via behavior-equivalent edits, and a single universal, carrier-constrained string optimized on a surrogate can transfer unchanged to black-box API targets. Sanitization can remove some insertion surfaces, but it rewrites many inputs and leaves executable-looking channels vulnerable. These findings collectively suggest that clean benchmark performance alone is insufficient as a security guarantee, and that integrity-oriented evaluations and defenses should explicitly target robustness to behavior-preserving code changes without resorting to overly conservative task-altering preprocessing.
Impact Statement
This work evaluates the robustness of LLM-based vulnerability detectors under semantics-preserving adversarial perturbations. We discuss both the positive contributions and potential risks of this research.
Positive Impacts. Our findings reveal essential blind spots in current LLM-based code security tools before they are exploited in the wild. By systematically characterizing failure modes and introducing metrics such as Complete Resistance, we provide actionable diagnostics for practitioners to harden detection pipelines. We believe responsible disclosure of these weaknesses, consistent with standard practices in security research, enables the community to develop more robust defenses.
Potential Risks and Mitigations. As with most security research, the techniques described could theoretically be misused to evade vulnerability detection. However, we note that: (1) the attack surfaces we study (identifier renaming, comment insertion, preprocessor directives) are already well-known code transformations; (2) we do not release optimized adversarial strings or attack toolkits; and (3) the primary barrier to exploiting vulnerabilities lies in discovering them, not in evading detection. We believe the defensive value of exposing these weaknesses substantially outweighs the marginal risk of informing attackers about evasion strategies they could likely discover independently.
Appendix A Detailed Threat Model
A.1 Target Models Details
We evaluate the proposed attacks across five state-of-the-art LLMs, categorized by their access regimes:
For reproducibility, we explicitly set temperature=0.0 for GPT-4o to ensure deterministic decoding, while GPT-5-mini was evaluated under its mandatory default settings, as the API does not currently support sampling parameter adjustment.
Regime Model Family Specific Version / API ID
White-box Llama 3.1 meta-llama/Llama-3.1-8B-Instruct
White-box CodeAstra rootxhacker/CodeAstra-7B (Ganesan, 2024)
White-box Qwen2.5-Coder Qwen/Qwen2.5-Coder-14B-Instruct (Surrogate)
Black-box (API) GPT-4o gpt-4o-2024-08-06
Black-box (API) GPT-5-mini gpt-5-mini-2025-08-07
Operational Black-box Qwen2.5-Coder Qwen/Qwen2.5-Coder-32B-Instruct
A.2 Vocabulary Constraints
To ensure that injected payloads do not break C/C++ syntax or preprocessor parsing, we restrict the GCG search space $\mathcal{A}$ based on the carrier type:
  • Identifiers and Dead-branch. Restricted to [A-Za-z_][A-Za-z0-9_]*.
  • Macro Bodies and Guards. We use a filtered printable ASCII set $\mathcal{A}_{\text{macro}}$ that excludes characters capable of terminating string literals or lines prematurely:
[…]
A.3 Semantics Preservation Validation
We validate semantics preservation through a two-tiered approach: (1) Constraint-based Validation to verify method-specific injection rules, and (2) Differential Syntactic Analysis using Tree-sitter to ensure that perturbations do not introduce new syntax errors compared to the original code. This approach is robust to the incomplete nature of function-level datasets (e.g., missing headers) while strictly flagging syntax violations.
Tier 1: Constraint-based Validation. For each attack method, we verify rigorous structural constraints:
  • Identifier substitution. (i) Replacement strings must match valid C/C++ identifier patterns ([A-Za-z_][A-Za-z0-9_]*); (ii) replacement must be consistent across the scope; (iii) Variable Shadowing Check. We verify that the new identifier does not conflict with existing variables in the local or global scope to prevent logic corruption.
  • Comment-carrier insertion. We verify that delimiters (/* */ or //) are balanced and boundaries are correctly recognized by a lexer.
  • Preprocessor-based insertion. We verify that directives (#ifdef/#endif) are strictly balanced and macro names use unique prefixes (e.g., SAFE_FUNC_) to avoid collisions.
  • Dead-branch code insertion. We verify that (i) the condition is statically false (e.g., if(0)), (ii) braces are balanced, and (iii) the block is syntactically distinct from surrounding code.
Tier 2: Differential Syntactic Analysis (Tree-sitter). Standard compilation fails on function snippets due to missing translation units (e.g., undefined types like u8). To address this, we employ Tree-sitter, a robust incremental parser. We define a transformation as Syntactically Preserved if the number of AST error nodes in the adversarial code ($E_{adv}$) does not exceed the error nodes in the original code ($E_{orig}$):
$\text{Validity}\iff E_{adv}\leq E_{orig}$
This differential metric filters out pre-existing errors caused by missing context (e.g., unknown macros) and isolates syntax errors introduced strictly by the attack.
Validation Results. Table 6 summarizes the validation rates across three diverse model families (Qwen, CodeAstra, and Llama). Constraint Satisfaction. Across all models, $99\%$ of samples satisfy method-specific constraints. Syntactic Validity. Under Tree-sitter differential analysis, our attacks achieve a consistent pass rate of 99.5% on average (e.g., 99.6% for Qwen2.5-Coder-32B and 99.5% for CodeAstra). Notably, the Tree-sitter Valid score for preprocessor-based carriers (#ifdef, Macro) and Comment carriers reaches 100%, while Identifier Substitution maintains $\approx 98.2\%$. This cross-model consistency confirms that our adversarial perturbations- whether applied to small or large test sets-preserve the syntactic structure of the original code, effectively ruling out malformed injections as a cause for evasion.
Table 6: Semantics Preservation Rates across Models. We report two metrics: (1) Constraint Sat. Fraction of samples satisfying method-specific rules. (2) Tree-sitter Valid. Fraction of samples where the attack introduces no new AST error nodes ($E_{adv}\leq E_{orig}$). High validity across diverse architectures confirms the stability of our attack generation.
Metric Id. Sub Comment #ifdef Macro Dead-Branch Overall
Model: Qwen2.5-Coder-32B ($N=1121$)
Constraint Sat. (%) 100.0 99.0 99.8 99.8 99.8 99.7
Tree-sitter Valid (%) 98.0 100.0 100.0 100.0 99.8 99.6
Model: CodeAstra ($N=3454$)
Constraint Sat. (%) 100.0 99.5 99.9 99.9 98.0 99.5
Tree-sitter Valid (%) 98.2 100.0 100.0 100.0 99.5 99.5
Model: Llama3.1-8B ($N=1552$)
Constraint Sat. (%) 100.0 99.4 99.8 99.8 97.9 99.4
Tree-sitter Valid (%) 98.6 100.0 100.0 100.0 99.4 99.6
Dataset Compilation Challenges. Function-level vulnerability datasets lack full compilation context (missing headers, type definitions, build dependencies). To enable rigorous compilation-based validation, we constructed a compilable subset from PrimeVul.
Tier 3: Compilation-based Validation. We constructed a Compilable Validation Set through:
  1. 1.
    Sample 2,000 ground-truth vulnerable functions from PrimeVul’s compilable subset
  2. 2.
    Verify all 2,000 compile standalone via gcc -fsyntax-only with standard headers
  3. 3.
    Evaluate on CodeAstra (highest TPR=69.08%) → 1,434 correctly detected (71.7%)
  4. 4.
    Apply transfer attacks using universal strings from Qwen2.5-Coder-14B surrogate
  5. 5.
    Test compilation preservation after attack
Header Injection Harness. stdio.h, stdlib.h, string.h, stdint.h, stddef.h, stdbool.h
Compiler flags: gcc -x c -fsyntax-only -Wno-implicit-function-declaration -
Validation Results. Table 7 shows compilation preservation rates. Additive carriers achieve perfect 100% preservation (1,434/1,434), providing compiler-verified proof of semantics-preservation by construction. Identifier Substitution achieves 90.9% (1,303/1,434). The 131 failures stem from collisions with injected header symbols (uint8_t, size_t, printf, malloc, etc.) and are excluded from ASR calculations.
Table 7: Compilation Preservation under Transfer Attacks (N=1,434N{=}1,434). CodeAstra true positives attacked with universal strings from Qwen-14B (Table 2). Validation via gcc -fsyntax-only.
Attack Carrier Baseline After Attack Preservation
Identifier Substitution 1,434 1,303 90.9%
Comment Insertion 1,434 1,434 100.0%
#ifdef Guard 1,434 1,434 100.0%
Macro Definition 1,434 1,434 100.0%
Dead-Branch 1,434 1,434 100.0%
Relationship to Main Results. Main evaluation (§5) uses full dataset (N=5,000, all models) with Tree-sitter validation (Tier 2). Tier 3 provides compiler-verified confirmation on the CodeAstra subset using identical transfer strings. The two tiers are complementary: Tier 2 scales to full evaluation; Tier 3 provides enough proof on a representative subset.
Limitations. Tier 3 covers CodeAstra predictions only, though transfer string universality means results generalize to other models. PrimeVul functions that require project-specific headers (28.3% of the original compilable sample) are excluded.
Appendix B Detailed Experimental Setup
B.1 Dataset Construction Details
Data sources and target language. Unified-VUL-N contains vulnerable functions written in C/C++ and isconstructed by pooling the vulnerable subsets from PrimeVul, BigVul, and DiverseVul.The goal is to evaluate robustness under semantics-preserving transformations on asingle, fixed benchmark with controlled composition.
Normalization for cross-source deduplication. For the sole purpose of cross-source deduplication, we build a normalized string code_norm for each function. The normalization is only for the purpose of deduplication, so the selected dataset still has comments. The normalization removes surface-level variation while preserving the underlying C/C++ token sequence:
  • Remove C-style block comments (/* ... */) and line comments (// ...).
  • Remove #-style comments that are not preprocessor directives (i.e., do not begin with a valid directive keyword), so that purely comment-like # lines do not affect deduplication.
  • Canonicalize newlines and collapse runs of whitespace (spaces/tabs) to a single space, without altering non-whitespace characters.
This normalization is not applied to the code presented to the model; it is only used to detect duplicates across datasets.
Deduplication rule. We compute $\texttt{code\_hash}=\texttt{SHA256}(\texttt{code\_norm})$ and treat two functions as duplicates if their hashes match. Deduplication is performed globally across the pooled data (i.e., across all three sources), ensuring that the final set contains no cross-source duplicates under the normalization above.
Context budget and token-length filtering. To guarantee sufficient headroom for injected strings and to standardize the context budget across models, we filter functions by code-only token length using the Qwen2.5-Coder-14B tokenizer. Specifically, we require the code-only length to be at most 4096 tokens, where “code-only” refers to the function text used as input code (excluding any additional prompt or JSON schema wrappers). This constraint ensures that our injected variants remain within model context limits under the shared evaluation prompt.
Quota-controlled deterministic sampling. After global deduplication and length filtering, we enforce fixed per-source quotas of 3000 (PrimeVul), 1000 (BigVul), and 1000 (DiverseVul). To make selection deterministic, we shuffle items within each source once using a fixed random seed (seed=42) and then select items in a fixed priority order (PrimeVul $\rightarrow$ BigVul $\rightarrow$ DiverseVul), skipping any item whose code_hash has already been selected. We continue until all quotas are met.
Evaluation uses original code (no normalization at inference). All evaluations (clean and attacked) use the original, unnormalized code for each function as provided by the source dataset, including comments and preprocessor directives. The normalized string is used only to define code_hash for deduplication.
Released artifacts. We release three artifacts to reproduce the dataset and enable audits:
  • JSONL. the final Unified-VUL-N set, with one record per function including the original function text and metadata (source, id, CWE label).
  • CSV manifest. a per-item manifest with (at minimum) source, id, cwe, func (original code), code_norm (dedup-normalized string), and code_hash (SHA-256 of code_norm).
  • Meta JSON. dataset-level metadata, including the random seed, tokenizer name, token budget, counts before/after normalization and deduplication, and the exact quota configuration.
B.2 Additional metric variants.
For a single attack family $A_{k}$, we also report
$\mathrm{ASR}_{\mathrm{cond}}(A_{k})=\frac{|\mathrm{Flip}(A_{k})|}{|\mathrm{TP}_{\mathrm{clean}}|},\qquad\mathrm{CR}(A_{k})=\frac{|\mathrm{Resist}(A_{k})|}{|\mathrm{TP}_{\mathrm{clean}}|}.$
Define the clean true-positive rate as $\mathrm{TPR}_{\mathrm{clean}}=\frac{|\mathrm{TP}_{\mathrm{clean}}|}{N}$. Under the union of attacks, the attacked true-positive rate is
$\mathrm{TPR}_{\mathrm{att}}=\frac{|\mathrm{TP}_{\mathrm{clean}}|-|\mathrm{Flip}_{\cup}|}{N},$
So the recall drop satisfies the equivalent form.
$\Delta\mathrm{TPR}=\mathrm{TPR}_{\mathrm{clean}}-\mathrm{TPR}_{\mathrm{att}}=\frac{|\mathrm{Flip}_{\cup}|}{N}.$
B.3 Benign-Class Invariance and False Positives
Setup. We evaluate benign-class invariance on a held-out set of $N_{\mathrm{benign}}{=}500$ ground-truth benign functions. Clean FPR is the fraction of benign functions predicted as VULNERABLE under clean inference (baseline model error). Sanitized FPR is computed after applying the same sanitization pipeline as in §5.3 (comment stripping and preprocessor removal); we report $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ as the change in false-positive rate (percentage points) from clean to sanitized inputs. Benign-class Invariance Rate is the fraction of benign functions that are predicted as BENIGN under clean inference and remain BENIGN after applying our semantics-preserving attacks (union over all attack methods). This metric directly measures whether our attacks preserve benign-class predictions, which is aligned with our attack objective (evading vulnerability detection without introducing false positives).
Results. Table 8 reports per-model clean FPR, sanitized FPR (and $\Delta\mathrm{FPR}_{\mathrm{sanit}}$), and benign-class invariance rate. All models achieve 100% benign-class invariance rate, demonstrating that our semantics-preserving attacks preserve benign-class predictions: all functions that are correctly classified as BENIGN under clean inference remain BENIGN after attack.
Table 8: Benign-side effects and invariance. Clean/Sanitized FPRs are measured on $N_{\mathrm{benign}}$ ground-truth benign functions. $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ is the change (percentage points) after sanitization. Benign-class Invariance is computed on benign functions predicted as BENIGN under clean inference and remains BENIGN after applying the union of semantics-preserving attacks.
Model $N_{\mathrm{benign}}$ Clean FPR (%) Sanitized FPR (%) $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ ($p$) Benign Invariance (%)
Qwen2.5-Coder-32B 500 11.0 (55/500) 9.6 (48/500) -1.4 100.0 (445/445)
GPT-4o 500 36.2 (181/500) 29.8 (149/500) -6.4 100.0 (319/319)
CodeAstra 500 70.0 (350/500) 63.4 (317/500) -6.6 100.0 (150/150)
Llama3.1-8B 500 16.8 (84/500) 18.0 (90/500) +1.2 100.0 (416/416)
GPT-5-mini 500 54.4 (272/500) 56.2 (281/500) +1.8 100.0 (228/228)
B.4 Robustness of Reference Tokenizer for Length Filtering
To ensure fair cross-model comparisons, we standardized length filtering using the Qwen2.5-Coder-14B tokenizer to maintain an identical evaluation set across all targets. Per-model filtering would result in inconsistent subsets, making metrics like $TPR_{clean}$ and $ASR_{cond}$ non-comparable. Correlation analysis on 500 random samples confirms that the Qwen tokenizer is highly representative: it achieves a Pearson correlation of $r=0.9998$ with Llama-3.1-8B and $r=0.9941$ with Mistral-7B-Instruct-v0.2, which is the original model of CodeAstra. The mean length difference relative to Llama is only 5.57-5.57 tokens (max 16). Given our 4096-token limit, these negligible deviations ensure that the filtering process does not introduce systemic bias against specific architectures.
B.5 Compute and Latency Statistics
We report compute and latency statistics for universal GCG optimization to assess practical attack cost. For the on-target identifier substitution experiment on Qwen2.5-Coder-32B (reported in §5, Transferability), we instrumented the GCG optimization loop to track forward passes and wall-clock time. The optimization uses curriculum learning where $m_{c}$ increases from 1 to $m_{\mathrm{train}}{=}10$ over the first 10 steps (one step per training function), then optimizes all $m_{\mathrm{train}}{=}10$ training prompts together for up to S=500S{=}500 total steps. Note that the optimization is performed only on the training set ($m_{\mathrm{train}}{=}10$ functions), not on the full evaluation set ($|\mathbf{TP}_{\mathrm{clean}}|{=}1121$ functions).
Optimization Hyperparameters. To ensure reproducibility, we used a fixed configuration for the GCG algorithm across all experiments. We set the optimization steps S=500S{=}500, candidate search width to $256$ (number of candidates generated per step), and top-$k$ filtering to $128$ for gradient guidance. Candidate evaluation used a batch size of $8$. The adversarial string $\sigma$ was initialized with a fixed placeholder pattern "x_x_x_x_x_x_x_x_x_x_" (length 20) to constrain the search space, and the optimization process used a fixed random seed of $42$. The support set $\mathcal{X}_{\text{opt}}$ (m=10m{=}10) was sampled randomly from the training split using a fixed seed of 21 to ensure reproducibility.
Hardware and software setup. We run experiments on NVIDIA H100 GPUs (typically 1 GPU per run). Models are loaded in bfloat16 precision without quantization. The tokenizer is Qwen2.5-Coder-14B HuggingFace tokenizer for length filtering; optimization uses the target model’s tokenizer (Qwen2.5-Coder-32B for on-target experiments).
Cost metrics. We track (i) forward passes for gradients ($m_{c}$ per step during curriculum, $m_{\mathrm{train}}$ per step after balance), (ii) forward passes for candidate evaluation ($m_{c}\cdot\lceil\texttt{search\_width}/\texttt{batch\_size}\rceil=m_{c}\cdot 8$ per step with our default settings), and (iii) wall-clock time (total optimization time including GPU memory transfers and batched operations). Optional per-step model.generate() calls are counted separately and can be disabled via enable_step_generate=False to reduce latency.
Black-box Query Cost Efficiency. An advantage of our transfer-based approach over query-based black-box attacks (e.g., PAIR, TAP, or genetic baselines) is the marginal cost of exploitation. Query-based methods typically require thousands of API interactions to optimize a single adversarial example for a specific target. In contrast, our method optimizes the universal string locally on a surrogate model (incurring zero API cost) and applies it to the target API using a single inference pass. For a target dataset of size $N$, query-based methods incur a cost of $O(N\times Q)$ where $Q$ is the query budget (often >1000>1000), whereas our approach incurs a cost of $O(N)$ (specifically, $1\times N$ queries). This makes our attack highly scalable and economically viable against commercial APIs like GPT-4o.
Per-method and per-model statistics. Detailed forward-pass counts and wall-clock times per attack method and model are available in our code repository. For the Qwen2.5-Coder-32B on-target identifier substitution run, the total optimization (including curriculum phase) completes in approximately 20 hours on a single H100. We note that universal GCG amortizes the optimization cost across all target functions, making it more efficient than per-instance optimization when attacking large evaluation sets.
Appendix C Supplementary Result
C.1 Site Selection and Functional Role
To investigate the sensitivity of LLM-based detectors to the spatial placement and syntactic role of adversarial strings, we conducted an ablation study on three models: Qwen2.5-Coder-32B, Llama3.1-8B, and CodeAstra. We focused on two primary dimensions: (i) the position of comment-carrier insertions (Leading, Middle, or Trailing) and (ii) the functional role of the adversarial pattern in macros (Name vs. Body).
Positional Sensitivity. Table 9 compares the $ASR_{cond}$ for different comment-carrier insertion sites. We observe that Trailing comments (placed after the function body) often exhibit high effectiveness, notably for Llama3.1-8B and CodeAstra, where success rates significantly exceed those of the original Leading header comments. However, Middle placements (interleaved with the logic) often show lower effectiveness across all models. This suggests that the model’s contextual focus on interleaved logic may partially mitigate the influence of semantics-preserving perturbations compared to injections at the function boundaries.
Functional Role Sensitivity. We further analyzed whether the specific role of a string-as part of a macro identifier (Name) versus a definition value (Body)-affects evasion success. Our results show that Macro Name attacks are considerably more potent across all targets. For example, CodeAstra’s resistance increases from $1.13\%$ under Name-based attacks to $69.22\%$ when the pattern is moved to the Body. This indicates that current detectors are particularly susceptible to adversarial signals when they are presented as active code identifiers rather than passive data.
Table 9: Ablation results for different injection sites and functional roles. $ASR_{cond}$ is reported as the fraction of originally correctly detected vulnerable functions ($TP_{clean}$) that were flipped to BENIGN.
Model Family Variant Flipped / $TP_{clean}$ $ASR_{cond}$ (%)
Qwen2.5-Coder-32B Comment Head (Original) 1026 / 1121 91.53
Middle 778 / 1121 69.40
Trailing 952 / 1121 84.92
Macro Name (Original) 962 / 1121 85.81
Body 621 / 1121 55.40
Llama3.1-8B Comment Head (Original) 652 / 1552 42.01
Middle 355 / 1552 22.87
Trailing 1163 / 1552 74.94
Macro Name (Original) 1016 / 1552 65.46
Body 595 / 1552 38.34
CodeAstra Comment Head (Original) 2178 / 3454 63.06
Middle 1565 / 3454 45.31
Trailing 2941 / 3454 85.15
Macro Name (Original) 3415 / 3454 98.87
Body 1063 / 3454 30.78
Appendix D Extended Sanitization Analysis
Table 10 details prediction drift when applying sanitization (stripping comments and preprocessor directives) to clean, unperturbed vulnerable functions. We measure $\Delta\mathrm{TP}=\mathrm{TP}_{\mathrm{sanitized}}-\mathrm{TP}_{\mathrm{original}}$, the change in true positive count.
Table 10: Sanitization-induced prediction drift. Stripping benign comments and preprocessor directives from clean code increases vulnerability detection, indicating that non-executable tokens act as lexical distractors.
[…]
Qwen2.5-Coder-32B 1,121 1,179 +58 (+5.17%)
Llama3.1-8B 1,552 1,685 +133 (+8.57%)
CodeAstra 3,454 3,457 +3 (+0.09%)
As shown in Table 10, the detection boundary is highly unstable. For Llama3.1-8B, simply removing developer comments (which have no effect on program execution) causes the model to ”discover” 133 previously missed vulnerabilities—an 8.6% increase. This suggests that the model’s clean accuracy is partly contingent on specific lexical distributions rather than robust semantic understanding. Qwen2.5-Coder-32B shows similar behavior (+5.2%), while CodeAstra exhibits minimal drift (+0.09%), suggesting different sensitivities to non-executable context across architectures.
This instability has two implications. First, it confirms that models use syntactically irrelevant tokens as decision features, violating semantic invariance. Second, it undermines sanitization as a defense: while it eliminates carrier surfaces, it simultaneously alters the model’s decision boundary in unpredictable ways, potentially introducing new false negatives or positives on benign code variations.
D.1 Code Length Effect: Additional Results
We measure length using whitespace-delimited token count, i.e., len(code.split()), and compute correlations with $\texttt{num\_successful\_methods}\in\{0,1,2,3,4,5\}$ (the number of carriers that flip a clean true positive) on the subset with complete outcomes across all five carriers.
Table 11: Correlation between code length and joint attack success. Pearson $r$ and Spearman $\rho$ are computed between whitespace token length and num_successful_methods.
Model Pearson $r$ Spearman $\rho$
GPT-4o 0.122-0.122 0.163-0.163
GPT-5-mini 0.172-0.172 0.208-0.208
Qwen2.5-Coder-32B 0.074-0.074 0.069-0.069
CodeAstra 0.068-0.068 +0.025+0.025
Llama3.1-8B 0.024-0.024 0.061-0.061
Interpretation. Across models, correlations are often negative but small in magnitude (Table 11), suggesting that longer functions tend to admit fewer successful carriers, with substantial variation by model. One plausible explanation is a dilution effect: as more semantic evidence is present in longer contexts, the marginal influence of a fixed, localized carrier (e.g., a header comment or macro name) may diminish. The binned curves in Figure 5 further indicate that the relationship is not purely linear.
Ceiling and variance effects. For Qwen2.5-Coder-32B (and to a lesser extent Llama3.1-8B), the length & robustness association is modest. The joint outcome num_successful_methods is skewed toward higher counts when multiple carriers succeed on many samples, which partially attenuates correlation-based summaries. Figure 5 visualizes these partial ceiling effects directly.
Carrier-specific heterogeneity. Length sensitivity is not uniform across carriers. At the carrier level, for Llama3.1-8B, macro-name carriers exhibit a clearer negative trend (Pearson $r\approx-0.163$; Spearman $\rho\approx-0.228$ for whitespace length vs. macro success), while other carriers show weaker dependencies. This pattern is suggestive of carrier-specific coupling, e.g., identifier-like roles may be more susceptible to dilution in longer contexts, but we leave controlled tests (e.g., fixing insertion position and matching structural complexity) to future work.
D.2 Temperature Robustness
Motivation. To verify that attack effectiveness is not an artifact of decoding stochasticity, we repeat the evaluation under two decoding temperatures, T=0T{=}0 (deterministic) and T=0.7T{=}0.7.
Protocol. For each target model and attack method, we run the full evaluation pipeline at both temperatures under the same prompts and parsing protocol. For stochastic decoding (T=0.7T{=}0.7), we report mean $\pm$ std over K=3K{=}3 runs.
Results. Table 12 shows that increasing decoding stochasticity does not qualitatively change susceptibility: the union success rate changes only slightly, and per-method success rates remain close across temperatures.
[…]
Attack Method T=0.0T{=}0.0 T=0.7T{=}0.7 (mean $\pm$ std)
ASR (union) 87.00% 85.19% $\pm$ 0.65%
gcg_var 31.09% 27.85% $\pm$ 0.68%
comment 65.48% 66.06% $\pm$ 0.99%
macro 42.35% 40.62% $\pm$ 0.92%
ifdef 71.61% 74.17% $\pm$ 0.64%
Appendix E Prompt templates
We used the following prompt template (implemented in code as ’get_prompt_template()’):
Strictly analyze code for specific vulnerable variables.
If the label is benign, use empty string for variable.
Output JSON only:
{"label":"VULNERABLE" or "BENIGN","variable":"<variable_name>"}
Do not output explanations. Stop after closing brace.
Code:
{func}
JSON:
Appendix F Attack Carrier Templates
To ensure full reproducibility, we provide the exact code templates used for each carrier family, corresponding to the implementation logic in our evaluation pipeline. In the templates below, [TRIGGER] denotes the GCG-optimized adversarial string (referred to as {optim_str} in the codebase), and [ORIGINAL_CODE] denotes the victim function text.
F.1 Comment Carrier
We inject a Doxygen-style multi-line comment block immediately preceding the function definition. The adversarial string is embedded within the @commit_hash metadata field, acting as a ”distractor” in the non-executable header.
 * @brief Function implementation
 * @commit_hash [TRIGGER]
 * @status verified
[ORIGINAL_CODE]
F.2 Macro Insertion
We prepend a macro definition before the function. The adversarial string is embedded into the macro identifier itself (e.g., SAFE_FUNC_...). This tests the model’s sensitivity to token sequences in the global scope that resemble function prototypes but are technically preprocessor directives.
#define SAFE_FUNC_[TRIGGER] benign_function()
[ORIGINAL_CODE]
F.3 Dead-Branch Insertion
Unlike the preprocessor carriers, this payload is injected inside the function body. We locate the final closing brace (}) of the function and insert a strictly dead branch (guarded by if(0)) immediately before it. The adversarial string is encapsulated within a benign string literal statement inside the block.
// ... inside the function body ...
    // [Insertion Point: before the final return/brace]
    if (0) {
        "benign_[TRIGGER]";
F.4 #ifdef Insertion
We wrap a benign log statement containing the adversarial string within an #ifdef DEBUG_MODE block. This payload is prepended to the function context, simulating a global or static debug configuration common in C/C++ projects. We verified that none of the 5,000 functions in our dataset define or reference common debug macros DEBUG_MODE, eliminating potential risk within our evaluation scope.
#ifdef DEBUG_MODE
    benign_log("[TRIGGER]");
#endif
[ORIGINAL_CODE]
F.5 Identifier Substitution
For the identifier substitution attack (not template-based but replacement-based), we identify the target variable $v$ (via the model’s "variable" attribution or frequency analysis) and apply a global string replacement:
$\texttt{code.replace}(v,\texttt{[TRIGGER]})$
We strictly validate that [TRIGGER] is a valid C identifier and does not shadow existing symbols in the local or global scope.
Figure 4: CWE $\times$ attack method success heatmaps across models. Each cell gives the per-method conditional attack success rate $\mathrm{ASR}_{\mathrm{cond}}(A;M)$ for a given CWE type and attack method. This visualization highlights strong method & CWE interactions and model-specific attack profiles (e.g., CodeAstra’s distinct behavior relative to other models).
Figure 5: Attack success rate versus code length (binned). Curves show per-carrier success rates across length ranges, complementing Table 11 by revealing non-linear trends and saturation regimes.

### v2

@@
-Semantics-Preserving Evasion of LLM Vulnerability Detectors
+Syntax- and Compilation-Preserving Evasion of LLM Vulnerability Detectors
 Abstract
-LLM-based vulnerability detectors are increasingly deployed in security-critical code review, yet their resilience to evasion under behavior-preserving edits remains poorly understood.
-We evaluate detection-time integrity under a semantics-preserving threat model by instantiating diverse behavior-preserving code transformations on a unified C/C++ benchmark (N=5000N{=}5000), and introduce a metric of joint robustness across different attack methods/carriers.
-Across models, we observe a systemic failure of semantic invariant adversarial transformations: even state-of-the-art vulnerability detectors perform well on clean inputs while predictions flip under behavior-equivalent edits.
-Universal adversarial strings optimized on a single surrogate model remain effective when transferred to black-box APIs, and gradient access can further amplify evasion success.
-These results show that even high-performing detectors are vulnerable to low-cost, semantics-preserving evasion.
-Our carrier-based metrics provide practical diagnostics for evaluating LLM-based code detectors.
-Machine Learning, ICML
+LLM-based vulnerability detectors are increasingly deployed in CI/CD security gating, yet their resilience to evasion under syntax- and compilation-preserving edits remains poorly understood.
+We evaluate five attack variants spanning four carrier families of behavior-preserving code transformations on a unified C/C++ benchmark (N=5000N{=}5000) and introduce Complete Resistance (CR), measuring the fraction of correctly detected vulnerabilities that withstand all attack variants.
+Our findings reveal a significant robustness gap: models achieving high clean recall exhibit CR as low as 0.12%, meaning over 87% of detected vulnerabilities can be evaded by at least one syntax-preserving edit.
+Universal adversarial strings optimized on a 14B surrogate transfer effectively to black-box APIs including GPT-4o, while on-target optimization further amplifies evasion (up to 99.88% ASR).
+These results indicate that clean benchmark accuracy alone is insufficient as a security guarantee for deployed vulnerability detectors.
+The project is available at https://anonymous.4open.science/r/code_review-6932.
 Figure 1: Why carrier-constrained optimization matters.
-Left: Carrier-constrained GCG via identifier substitution preserves compilation and flips the detector to BENIGN.
-Middle: Naively appending a GCG-optimized string breaks compilation, so it is not a valid semantics-preserving evasion.
-Right: Unoptimized semantics-preserving edits remain compilable but typically fail to evade detection.
-Red denotes the GCG-optimized identifier, blue denotes a random identifier, and boldface highlights the original identifier.
+Left (red=GCG-optimized, blue=random, bold=original): Carrier-constrained GCG via identifier substitution preserves compilation and flips the detector to BENIGN.
+Middle: Naively appending a GCG-optimized string breaks compilation, so it is not a valid syntax-preserving evasion.
+Right: Unoptimized syntax-preserving edits remain compilable but typically fail to evade detection.
 1 Introduction
 Software vulnerabilities remain among the most exploited attack vectors, driving rapid adoption of automated detection in modern continuous integration pipelines (Verizon, 2025; Microsoft, 2025).
-LLM-based detectors are now integrated into code review workflows at scale.
-This raises a security-critical question: do such detectors satisfy semantic invariance, or specifically, can an attacker evade detection using perturbations where the transformed program remains strictly compilable and retains the underlying vulnerability?
-If the latter holds, vulnerable code could reliably slip through automated security checks via innocuous transformations, including renaming identifiers, inserting comments, injecting dead-branch code, or adding preprocessor-guarded code, while leaving the underlying exploit fully intact.
-Existing research has highlighted the broad spectrum of adversarial vulnerabilities in LLMs (Fu et al., 2024; Vassilev et al., 2025).
+LLM-based detectors are now integrated into CI/CD gating workflows at scale, functioning as automated pass/fail checkpoints before code is merged.
+This raises a security-critical question: do such detectors satisfy semantic invariance, or specifically, can an attacker evade detection using perturbations where the transformed program remains strictly compilable and leaves the underlying vulnerability intact at the source level?
+If the latter holds, vulnerable code could reliably slip through automated security gates via innocuous transformations such as renaming identifiers, inserting comments, injecting dead-branch code, or adding preprocessor-guarded blocks.
+Existing research has highlighted various adversarial vulnerabilities in LLMs (Fu et al., 2024; Vassilev et al., 2025).
 Prior adversarial methods often ignore syntactic constraints, producing perturbations that break compilation or alter program logic (Ebrahimi et al., 2018).
 These invalid edits are practically useless: they cannot preserve the vulnerability’s executable behavior.
-As illustrated in Figure 1, effective evasion presents a dual challenge: naive GCG attack (Middle) breaks compilation, while randomized semantic edits (Right) often fail to evade detection.
+As illustrated in Figure 1, effective evasion presents a dual challenge: naive gradient-based attack (Middle) breaks compilation, while randomized syntax-preserving edits (Right) often fail to evade detection.
 A valid attack must simultaneously satisfy the compiler’s strict syntax rules and the adversary’s evasion objective (Left).
 More general-purpose LLM safety evaluations, on the other hand, primarily prioritize preventing harmful outputs, such as toxic content and jailbreak-style policy violations (Derczynski et al., 2024; Wei et al., 2025).
-Vulnerability detection, however, confronts an orthogonal integrity threat: adversaries need not induce malicious generation; they only need to make the detector oblivious to an existing vulnerability through behavior-preserving edits.
-This exposes a gap in current evaluation practices: standard safety probes emphasize output harmfulness, while clean-accuracy benchmarks ignore robustness under semantic-preserving transformations.
+Vulnerability detection, however, confronts an orthogonal integrity threat: adversaries need not induce malicious generation; they only need to make the detector oblivious to an existing vulnerability through syntax-preserving edits.
+This exposes a gap in current evaluation practices: standard safety probes emphasize output harmfulness, while clean-accuracy benchmarks ignore robustness under syntax-preserving transformations.
 Consequently, a detector can appear safe under toxicity-focused evaluations and achieve strong benchmark scores, yet remain brittle to innocuous perturbations in real-world deployment.
-To bridge this gap, we present a systematic evaluation framework for detection-time integrity under a semantics-preserving threat model.
-We instantiate multiple behavior-equivalent attack surfaces and evaluate both open-weight and proprietary models (via surrogate transfer) on a unified C/C++ benchmark.
-Beyond per-attack evasion, we introduce complete resistance, measuring the fraction of vulnerabilities that withstand all semantics-preserving transformations.
+We address this gap with a systematic evaluation framework for detection-time integrity under a syntax-preserving threat model.
+We instantiate multiple syntax-preserving attack surfaces and evaluate both open-weight and proprietary models (via surrogate transfer) on a unified C/C++ benchmark.
+Beyond per-attack evasion, we introduce complete resistance, measuring the fraction of vulnerabilities that withstand all five evaluated attack variants simultaneously.
 We make the following contributions:
   1. 1.
-    First Gradient-Based Evaluation of Semantic Invariance under Carrier Constraints.
-We combine gradient-driven optimization with carrier-constrained perturbations to expose LLM detector vulnerabilities.
-Preprocessor-style carriers (#ifdef, macros) achieve high success rates via universal triggers, contrasting with search-based methods (Jiang et al., 2026).
-We introduce Complete Resistance (CR) to measure joint robustness across carriers.
+    Carrier-Constrained Gradient-Based Evaluation of LLM Vulnerability Detectors.
+We combine gradient-based universal optimization with carrier-constrained perturbations to expose detection-time fragility.
+We introduce gradient-guided optimization with compiler-validated carrier constraints for LLM-based detectors, and propose Complete Resistance (CR) to measure joint robustness across all carriers simultaneously.
   2. 2.
     Exposing the Fragility of High-Performing Detectors.
-We show that models achieving 70%+ clean accuracy collapse under minimal perturbations: fewer than 13% of vulnerabilities resist all carrier types (CR << 13%), meaning 87% of ”detected” threats can be evaded by at least one semantics-preserving edit (e.g., #ifdef insertion).
-High benchmark performance masks critical robustness failures.
-Detectors succeed on benign inputs but fail severely when code undergoes innocuous transformations.
+Across five of six evaluated models, CR $\leq$ 13%: over 87% of correctly detected vulnerabilities can be evaded by at least one syntax-preserving edit (details in §5 and §5.4).
+High clean recall masks critical robustness failures.
   3. 3.
     High Exploitability of Black-box APIs via Low-barrier Transfer.
-We reveal that universal adversarial strings optimized on small surrogate models transfer effectively to closed-source models like GPT-4o without additional optimization.
+We show that universal adversarial strings optimized on small surrogate models transfer effectively to closed-source models like GPT-4o without additional optimization.
 2 Related Work
@@
 Early datasets like BigVul and DiverseVul contain label noise and train-test leakage (Chen et al., 2023; Fan et al., 2020), while PrimeVul addresses these issues through manual curation and stricter deduplication (Ding et al., 2024).
-Since our adversarial evaluation requires clean labels-a mislabeled ”vulnerable” sample would falsely appear robust-we adopt PrimeVul’s protocols and extend them with carrier-specific controls (§4.1).
+Since our adversarial evaluation requires clean labels a mislabeled "vulnerable" sample would falsely appear robust, we adopt PrimeVul’s protocols and extend them with carrier-specific controls (§4.1).
 LLMs for Vulnerability Detection.
-Recent benchmarks, such as (Khare et al., 2025), report that state-of-the-art LLMs can achieve promising performance on standard datasets.
-However, other studies indicate that code language models remain unreliable for detection in realistic settings, paralleling similar fragility observed in code reasoning tasks (Orvalho & Kwiatkowska, 2025), highlighting the need to evaluate robustness beyond clean accuracy (Ding et al., 2024).
-While separate efforts benchmark the security of generated code (e.g., SafeGen-style evaluations), our focus is strictly on detection-time robustness under semantics-preserving edits, not on generation safety (Li et al., 2025).
-Concurrently, recent work studies input-side defenses for adversarial transformations on code LMs.
-Evaluate-and-Purify proposes EP-Shield, using an LLM-as-a-judge to assess naturalness and purify identifier-substitution adversarial examples to restore predictions (Mu et al., 2025).
-Our work is complementary: we focus on attack-centric robustness evaluation under a semantics-preserving threat model spanning identifier substitution, comment-carrier insertion, preprocessor-based insertions, and dead-branch code carrier, and report metrics (ASR, Complete Resistance, Transferability) that can also be used to quantify the gains of such purification wrappers across attack surfaces.
+Recent benchmarks report that state-of-the-art LLMs can achieve strong performance on standard datasets (Khare et al., 2025), but code language models remain unreliable for detection in realistic settings (Orvalho and Kwiatkowska, 2025), highlighting the need to evaluate robustness beyond clean accuracy (Ding et al., 2024).
+Concurrently, Evaluate-and-Purify proposes EP-Shield, using an LLM-as-a-judge to purify identifier-substitution adversarial examples (Mu et al., 2025).
+Our work is complementary: we focus on attack-centric evaluation spanning four carrier families, and report metrics (attack success rate, complete resistance, transferability) that can quantify EP-Shield’s gains across attack surfaces.
 Code-Specific Adversarial Attacks on Vulnerability Detection.
-Adversarial attacks on code models span multiple representations.
-NatGVD (Rath et al., 2025) modifies graph structures for GNN detectors, while neural binary analyzers face syntactic mutation risks (Bundt et al., 2023).
-We introduce gradient-driven, carrier-constrained analysis to address this gap.
-Our universal triggers reveal that preprocessor carriers achieve 18%–99% ASR variance across models, fundamentally differing from identifier-based attacks.
-We propose Complete Resistance (CR) to measure joint robustness, a dimension orthogonal to SVulAttack’s instance-specific optimization.
+NatGVD (Rath et al., 2025) modifies graph structures for GNN detectors.
+Yefet et al.
+(2020) introduce adversarial identifier renaming and dead-code insertion for code models.
+HogVul (Yang et al., 2026) combines identifier and structural edits via PSO for pre-trained code models with compilation guarantees; our work differs in targeting instruction-tuned LLM detectors with gradient-based universal optimization, cross-regime evaluation including closed APIs, and a joint robustness metric (CR) across carriers.
+EaTVul (Liu et al., 2024) and Natural Attack (Yang et al., 2022) provide ChatGPT-generated and BERT-guided baselines without gradient optimization or compilation validation; we compare both via random identifier substitution (§5).
+We propose Complete Resistance (CR) to measure joint robustness across all carrier variants simultaneously.
 Jailbreaks and Transferable Prompt Attacks.
-Adversarial attacks on aligned LLMs have evolved from manual template engineering to automated frameworks.
-Optimization-based methods, such as GCG (Zou et al., 2023) and AutoDAN (Zhu et al., 2023), utilize gradient-guided or genetic search to find transferable adversarial suffixes.
-Concurrently, LLM-guided black-box attacks like PAIR (Chao et al., 2024), decomposition attack (Brown et al., 2025), and TAP (Mehrotra et al., 2024) leverage attacker models to iteratively refine prompts.
-Recent studies have extended this by characterizing the “utility” of such outputs (Nikolić et al., 2025).
-Existing evaluations, however, rarely examine how such transferable prompts interact with code tasks in a detection setting where edits must preserve compilability and semantics.
+Adversarial attacks on aligned LLMs range from manual template engineering to automated frameworks.
+Optimization-based methods, such as GCG (Zou et al., 2023) and AutoDAN (Zhu et al., 2023), use gradient-guided or genetic search to find transferable adversarial suffixes.
+Concurrently, LLM-guided black-box attacks like PAIR (Chao et al., 2024), decomposition attack (Brown et al., 2025), and TAP (Mehrotra et al., 2024) use attacker models to iteratively refine prompts.
+Recent studies have extended this by characterizing the "utility" of such outputs (nikolić2025jailbreaktaxusefuljailbreak).
+Existing evaluations, however, rarely examine how such transferable prompts interact with code tasks in a detection setting where edits must preserve compilability and syntactic structure.
 We adapt GCG to this constrained setting.
@@
 3.1 Threat Model
-Adversary objective.
-We study evasion attacks against LLM-based vulnerability detectors.
-Given a ground-truth VULNERABLE function $x$ and a detector $f(\cdot)$, the adversary aims to produce a perturbed function $x^{\prime}$ such that $f(x^{\prime})=\texttt{BENIGN}$ (a false negative), allowing the vulnerability to bypass automated security gates.
-Capabilities and semantic-invariance constraint.
-The adversary may modify only the source-code input while aiming to preserve semantic invariance; the transformed program must remain compilable and retain the underlying vulnerability.
-We restrict edits to four carriers that are intended to be semantics-preserving by construction: (i) scope-consistent identifier substitution, (ii) non-executable comment carriers, (iii) preprocessor carriers (macros/conditional guards) that are inactive under our compilation setting (e.g., unused macro definitions or disabled guards), and (iv) dead-branch insertion with a statically false predicate.
-We automatically validate carrier well-formedness (scope-consistent replacement, balanced comment delimiters, well-nested preprocessor directives/macro definitions, and syntactically valid dead-branch templates); we report pass rates and retain all samples without post-hoc filtering, validated by two-tier analysis: differential syntactic analysis (Tree-sitter (Khare et al., 2025)) on 5,000 functions and compilation-based validation (gcc) on another 1,434 standalone-compilable functions (Appendix A.3, Tables 67).
+Scenario.
+We study a CI/CD evasion scenario: an adversary controls existing vulnerable code that an automated LLM-based scanner flags as VULNERABLE and must pass the scan while keeping the vulnerability intact.
+This setting is relevant because LLM-based vulnerability scanners are embedded in CI/CD pipelines as gating mechanisms (Verizon, 2025; Microsoft, 2025), pass/fail checkpoints that operate without human review.
+The adversary does not inject a new vulnerability from benign code; they start from code that is already flagged and apply source-level edits that leave the exploit’s code structure unchanged.
+Carriers are non-interfering by construction (non-executable or unreachable at runtime), so the underlying vulnerability remains intact at the source level.
+Terminology.
+A carrier $c$ is a syntactic vehicle (identifier substitution, comment, preprocessor, dead-branch) for injecting adversarial string $\sigma$ such that compilation and runtime behavior are unaffected; we evaluate four carrier families yielding five attack variants (macro and #ifdef form the preprocessor family, each evaluated separately; §3.3).
+$\mathcal{X}_{\text{opt}}$ denotes the fixed support set of m=10m{=}10 diverse vulnerable functions used to optimize a universal $\sigma$ (GCG-optimized adversarial token sequence).
+$\mathcal{T}_{c}(x,\sigma)$ is the transformation operator that injects $\sigma$ into code $x$ via carrier $c$, producing adversarial variant $x^{\prime}$.
+We use syntax-preserving edits to denote transformations validated at the syntax level (Tree-sitter) and compilation level (gcc); formal runtime-equivalence is not asserted.
+Adversary.
+Given a ground-truth VULNERABLE function $x$ and detector $f(\cdot)$, the adversary produces $x^{\prime}=\mathcal{T}_{c}(x,\sigma)$ such that $f(x^{\prime})=\texttt{BENIGN}$.
+Edits are restricted to four carriers that are syntax-preserving by construction and largely compilation-preserving: (i) scope-consistent identifier substitution, (ii) non-executable comment carriers, (iii) preprocessor carriers (macros/#ifdef guards) inactive under our compilation setting, and (iv) dead-branch insertion with a statically false predicate.
+Carrier well-formedness is automatically validated; pass rates are reported in Appendix A.3.
+We evaluate this adversary under three realistic access assumptions that span the practical threat landscape.
 Knowledge regimes. We evaluate robustness under three realistic access assumptions:
@@
   • Black-box (API) via transfer.
-Targeting proprietary models with inaccessible internals, the adversary optimizes universal strings on a smaller surrogate model and applies them to the target without further modification.
-  • Operational black-box.
-Targeting large locally-hosted models where weights are accessible but direct gradient-based optimization is computationally expensive for typical adversaries.
-We primarily evaluate this regime using transfer-only methods to establish a realistic baseline of low-cost exploitability.
-However, to quantify the gap between this low-cost approach and the worst-case security risk, we explicitly perform an on-target white-box ablation (see Table 5) as a high-cost upper bound.
-Detailed model specifications and versions are provided in Appendix A.1.
+Targeting proprietary models with inaccessible internals, the adversary optimizes universal strings on a surrogate model and applies them to the target without modification.
+  • Cost-limited gray-box.
+Model weights are accessible but on-target gradient optimization is computationally prohibitive for most adversaries (e.g., single-H100 GCG on a 32B model takes $\sim$20 hours per carrier; §B.5).
+We evaluate this regime via transfer-only attacks as a practical lower bound on exploitability, and provide an on-target white-box ablation (Table 8, Appendix C.2) as a worst-case upper bound.
+Scope and validation boundary.
+This model covers automated detectors operating on uncompiled source code (CI pre-commit hooks, IDE integrations, API-based scanners).
+It does not address human code review or ensemble systems combining LLM reasoning with classical static/dynamic analysis.
+Cost comparison between this evasion route and introducing new undetectable code is deployment-specific and outside our scope.
+Full validation methodology (three-tiered: constraint-based checks for method-specific injection rules; Tree-sitter differential analysis on 5,000 functions; gcc compilation on a 1,434-function compilable subset) and preprocessor inactivity guarantees are detailed in Appendix A.3.
 3.2 GCG-based Optimization
@@
 We use the Greedy Coordinate Gradient (GCG) algorithm to optimize a discrete, printable adversarial string $\sigma$ that induces a prediction flip from VULNERABLE to BENIGN.
-GCG uses token-level gradients to guide coordinate-wise updates in a discrete search space, while all edits are restricted to the semantics-preserving transformation space defined by our threat model.
+GCG uses token-level gradients to guide coordinate-wise updates in a discrete search space, while all edits are restricted to the syntax-preserving transformation space defined by our threat model.
 We choose GCG for its ability to produce a universal string that transfers across both samples and models (§5.4).
 Optimization Objective and Universal Strategy.
-The attacker maximizes the model’s relative preference for the target label BENIGN over the ground truth VULNERABLE.
-Specifically, for a code instance $x$ and carrier $\mathcal{T}(x,\sigma)$, we maximize the log-odds objective $\mathcal{J}(\sigma;x)=\log p_{\theta}(\texttt{BENIGN}\mid\mathcal{T}(x,\sigma))-\log p_{\theta}(\texttt{VULNERABLE}\mid\mathcal{T}(x,\sigma))$.
-To ensure generalization, rather than optimizing per-instance, we learn a single universal adversarial string $\sigma$ over a fixed support set $\mathcal{X}_{\text{opt}}$ (m=10m{=}10) by maximizing the mean performance: $\mathcal{J}_{\text{univ}}(\sigma)=\frac{1}{m}\sum_{x\in\mathcal{X}_{\text{opt}}}\mathcal{J}(\sigma;x)$.
-Post-optimization, $\sigma$ is frozen to evaluate zero-shot transferability.
-The success of such input-agnostic triggers suggests that the detector’s vulnerability stems from systematic feature bias rather than instance-specific brittle directions.
-Implementation details. Optimization hyperparameters, surrogate model selection (Qwen2.5-Coder-14B), and prompt templates are provided in Appendix B.5 and Appendix E.
+The attacker maximizes the preference of the model for the target label BENIGN over the ground truth VULNERABLE.
+Specifically, for a code instance $x$ and carrier $c$, we maximize the log-odds objective $\mathcal{J}(\sigma;x)=\log p_{\theta}(\texttt{BENIGN}\mid\mathcal{T}_{c}(x,\sigma))-\log p_{\theta}(\texttt{VULNERABLE}\mid\mathcal{T}_{c}(x,\sigma))$.
+To ensure generalization, rather than optimizing per-instance, we learn a universal adversarial string $\sigma$ over the fixed support set $\mathcal{X}_{\text{opt}}$ (m=10m{=}10) by maximizing the mean objective: $\mathcal{J}_{\text{univ}}(\sigma)=\frac{1}{m}\sum_{x\in\mathcal{X}_{\text{opt}}}\mathcal{J}(\sigma;x)$.
+Once optimized, $\sigma$ is frozen to evaluate zero-shot transferability.
+This construction parallels universal adversarial triggers for NLP (Wallace et al., 2021), adapted to a carrier-constrained code domain with compiler-validated syntax and compilation preservation (Tree-sitter on 5,000 functions; gcc on 1,434-function subset).
+The success of such triggers suggests that the detectors rely on systematic feature bias rather than sample-level artifacts.
+Implementation details. Optimization hyperparameters, surrogate model selection (Qwen2.5-Coder-14B), and prompt templates are provided in Appendix B.5 and Appendix F.
 3.3 Attack Categories
-We define the transformation operator $\mathcal{T}(x,\sigma)$ through four families of insertion carriers.
-The following methods ensure functional equivalence by construction, modifying only the source-code input while strictly maintaining compilability and executable behavior (see Appendix F for details):
-  1. 1.
+We define the transformation operator $\mathcal{T}_{c}(x,\sigma)$ through four carrier families, yielding five attack variants evaluated in all tables (identifier substitution, comment, macro, #ifdef, dead-branch).
+The preprocessor family contains two variants (macro and #ifdef) sharing alphabet constraints but applied at distinct syntactic sites.
+Each carrier ensures syntax and compilation preservation by construction (see Appendix G for templates; carrier-specific alphabet constraints in Appendix A.2):
 Identifier substitution.
-We apply a consistent renaming map $\mathcal{M}:v\mapsto\sigma$ to the specific identifier $v$ explicitly flagged by the detector (i.e., extracted from the "variable" JSON field), or to the most frequent local identifier if attribution is unavailable.
-We replace all occurrences of $v$ within its scope.
-This focuses the perturbation on lexical cues without affecting variable binding or logic (avoiding variable shadowing).
-  2. 2.
+We apply a consistent renaming map $\mathcal{M}:v\mapsto\sigma$ to the detector-flagged identifier $v$ (or the most frequent local identifier if attribution is unavailable), replacing all occurrences within its scope without affecting variable binding (no shadowing; implementation in Appendix G).
 Comment-carrier insertion.
@@
 Comments act as transparent containers for adversarial tokens, allowing the insertion of non-interpretable strings.
-  3. 3.
-    Preprocessor-based insertion (macro/guard carriers).
-We inject $\sigma$ into inactive preprocessor regions, specifically unused macro definitions (denoted macro_name) or disabled conditional guards.
-These regions are stripped or ignored during compilation, ensuring zero impact on runtime behavior.
-  4. 4.
+[…]
 Dead-branch code insertion.
-We inject $\sigma$ into a formally unreachable control-flow block (e.g., if (0) { }).
+We inject $\sigma$ into an unreachable control-flow block (e.g., if (0) { }).
 Unlike preprocessor directives, this carrier is syntactically indistinguishable from executable code, serving as a diagnostic for the detector’s ability to resolve basic Boolean control flow.
-Family-specific candidate constraints.
-To ensure GCG-optimized strings remain valid under each carrier, we apply family-specific constraints to the candidate alphabet.
-For identifier substitution and dead-branch code, we enforce standard C-identifier compliance (alphanumeric with underscores, non-numeric start).
-For macro bodies and preprocessor-guarded blocks, we restrict the search space to a printable subset and enforce single-line constraints to prevent preprocessor syntax errors.
-Detailed vocabulary definitions and filtering rules are provided in Appendix A.2.
-Default carriers and ablations.
-Unless otherwise specified, comment-carrier insertion uses a leading/header carrier, and macro insertion uses the macro-name carrier (macro_name).
-We study alternative placements (middle/trailing) and the macro-body carrier (macro_body) only in a focused ablation (Table 9).
 4 Experimental Setup
 4.1 Datasets Construction
-Unified benchmark and design goals.
-To isolate robustness to semantics-preserving edits from confounders such as duplication leakage and length artifacts, we construct a single unified benchmark, Unified-VUL-N (N=5000N{=}5000).
-We pool vulnerable functions from PrimeVul, BigVul, and DiverseVul, enforcing three strict controls: (i) global cross-source de-duplication to prevent inflation from repeated functions, (ii) a fixed source composition to standardize evaluation across models, and (iii) a shared prompt budget to ensure valid comparisons.
-Deduplication by normalization-only hashing.
-To prevent evaluation inflation from surface-level duplicates, we canonicalize code logic (stripping comments and formatting) to define a deterministic equivalence class via SHA-256 hashing.
-This normalization is used strictly for cross-source auditing; models are always evaluated on the original, unmodified source code.
-Detailed normalization rules are provided in Appendix B.1.
-Filtering and Sampling.
-We restrict code to $\leq 4096$ tokens to reserve headroom for adversarial carriers and eliminate truncation-driven artifacts.
-A deterministic, quota-controlled sampling strategy ensures a stable and auditable benchmark composition (also see Appendix B.1).
+Unified benchmark.
+We construct Unified-VUL-N (N=5000N{=}5000) by pooling vulnerable C/C++ functions from PrimeVul, BigVul, and DiverseVul with three controls: global cross-source SHA-256 deduplication (on comment-stripped, whitespace-normalized code), a fixed per-source composition (PrimeVul 3000, BigVul 1000, DiverseVul 1000), and a shared prompt budget (${\leq}4096$ tokens via Qwen2.5-Coder-14B tokenizer).
+Deduplication targets cross-source inflation; models are always evaluated on the original, unmodified source code.
+Full normalization rules and sampling procedure are provided in Appendix B.1.
 4.2 Evaluation Protocol
@@
 Empirically, output stability is high ($99.9\%$ valid parse rate), confirming that measured failures stem from semantic understanding gaps, not syntax degradation.
-Evaluation Decomposition: Conditional Evasion.
-To isolate robustness from clean-accuracy capabilities, we evaluate attacks conditionally on Clean True Positives ($\mathrm{TP}_{\mathrm{clean}}$).
-Since a pre-existing error cannot be ”further evaded,” applying attacks to the entire dataset would inflate perceived robustness.
-We define $\mathrm{Flip}(A)$ as the subset of $\mathrm{TP}_{\mathrm{clean}}$ successfully evaded under attack $A$, and $\mathrm{Resist}(A)$ as those that maintain detection.
-Primary Robustness Metrics. We report three metrics to characterize the Robustness & Accuracy Gap:
+Conditional Evasion.
+We evaluate attacks on Clean True Positives ($\mathrm{TP}_{\mathrm{clean}}$) only: $\mathrm{Flip}(A)$ is the subset flipped under attack $A$; $\mathrm{Resist}(A)$ is the subset that maintains detection.
+Applying attacks to pre-existing misses would inflate apparent robustness.
+Robustness Metrics. We report three metrics to characterize the Robustness & Accuracy Gap:
   1. 1.
@@
   2. 2.
-    Complete Resistance (CR).
-The fraction of vulnerabilities that withstand all evaluated attack families ($\cap_{k}\mathrm{Resist}(A_{k})/|\mathrm{TP}_{\mathrm{clean}}|$).
-This serves as the security floor, indicating the proportion of reliable detections that are truly invariant to semantic perturbations.
+    Complete Resistance ($\mathrm{CR}$).
+The fraction of vulnerabilities that resist all evaluated attack variants: $\mathrm{CR}={\bigl|\bigcap_{k\in\mathcal{K}}\mathrm{Resist}(A_{k})\bigr|}\big/{|\mathrm{TP}_{\mathrm{clean}}|}$ where $\mathcal{K}=\{\text{idsub, comment, macro, {\#ifdef}, dead-branch}\}$.
+Each $A_{k}$ uses one canonical insertion site per variant (leading comment header, macro-name, #ifdef DEBUG_MODE, dead-branch before final brace; ablations over placements are in Table 6).
+CR is a lower bound on the security floor under the evaluated threat set: a vulnerability counted in CR resists every canonical attack; a vulnerability not in CR can be evaded by at least one.
+Because CR is defined relative to $\mathcal{K}$, it can only decrease as $\mathcal{K}$ is extended (e.g., adding more carrier positions or surrogate models).
   3. 3.
@@
 The absolute decrease in detection coverage on the full dataset ($|\mathrm{Flip}_{\cup}|/N$).
-This measures the operational risk of deploying the model in an adversarial environment.
+This measures the risk of deploying the model in an adversarial environment.
 Formal derivations and single-family variants are detailed in Appendix B.2.
+For random_idsub, we report mean $\pm$ std over K=3K{=}3 random seeds.
+GCG-based attacks use a fixed seed and support set $\mathcal{X}_{\text{opt}}$; point estimates reflect deterministic optimization under fixed conditions.
+Statistical qualification.
+Multi-seed analysis (on-target identifier substitution, N=3N{=}3 support sets, seeds 1/21/42) yields $\mathrm{ASR}_{\mathrm{cond}}$: CodeAstra $80.7\%{\pm}9.1$, Llama3.1-8B $31.1\%{\pm}11.3$, StarCoder2-15B $17.6\%{\pm}6.8$.
+All seeds produce successful attacks, confirming systematic exploitability; spread (${\leq}12$ pp) reflects support-set variability.
+Other model/carrier values are single-run estimates under seed 21 with fixed $\mathcal{X}_{\text{opt}}$.
 Baseline Clean Coverage. All attack metrics are contextualized against the clean performance baselines reported in Table 1, establishing the performance ceiling from which robustness degrades.
-Table 1: Baseline clean coverage on the unified C/C++ evaluation set.
-Absolute counts and $\mathrm{TPR}_{\mathrm{clean}}$ are reported; attack metrics ($|\mathbf{Flip}_{\mathrm{union}}|$, $|\mathbf{Resist}_{\mathrm{complete}}|$) appear in §5.1.
-Model NN $|\mathbf{TP}_{\mathrm{clean}}|$ $|\mathbf{FN}_{\mathrm{clean}}|$ $\mathrm{TPR}_{\mathrm{clean}}$ (%)
-Qwen2.5-Coder-32B 5000 1121 3879 22.42
-Llama3.1-8B 5000 1552 3448 31.04
-CodeAstra 5000 3454 1546 69.08
-GPT-4o 5000 2300 2700 46.00
-GPT-5-mini 5000 3681 1319 73.62
 5 Results
@@
 One would expect that a detector achieving 70% clean accuracy provides reliable security coverage for the vulnerabilities it identifies.
-Our evaluation reveals the opposite: across all models except GPT-5-mini, over 87% of correctly detected vulnerabilities can be flipped to false negatives using only semantics-preserving edits (Figure 2).
-CodeAstra, despite correctly flagging 3,454 vulnerabilities on clean inputs, maintains robust detection for only 4 functions under on-target attacks (CR = 0.12%), a 99.88% collapse in detection integrity.
-This pattern occurs under realistic access assumptions: surrogate-transfer attacks on black-box APIs (GPT-4o, GPT-5-mini) and operational black-box models (Qwen2.5-Coder-32B), supplemented by on-target optimization for accessible models (Llama3.1-8B, CodeAstra) to establish upper bounds.
-Across this threat landscape, GPT-5-mini is an exception, exhibiting substantially lower attack success (44.04% union $\mathrm{ASR}_{\mathrm{cond}}$) compared to other models (87%–100%), suggesting stronger invariance to carrier-based perturbations.
-These results indicate that clean benchmark accuracy is not a security guarantee.
-An attacker needs only to evade the subset of cases a detector would otherwise catch, and our results show this is achievable through innocuous transformations like identifier renaming or comment insertion (Table 2; clean baseline in Table 1).
-Figure 2: Overall attack effectiveness across models.
-Solid bars: transfer attacks using universal strings from Qwen2.5-Coder-14B surrogate.
-Hatched bars: on-target optimization (Llama3.1-8B, CodeAstra only).
-(a) Union $\mathrm{ASR}_{\mathrm{cond}}$ across carriers.
-(b) End-to-end recall decrease ($\Delta\mathrm{TPR}$).
+Our evaluation reveals the opposite: across all models except GPT-5-mini, over 87% of correctly detected vulnerabilities can be flipped to false negatives using only syntax-preserving edits (Tree-sitter-validated; compilation-preserving for comment, preprocessor, and dead-branch carriers and 90.9%-preserving for identifier substitution on the compilable subset; Table 1).
+CodeAstra, despite correctly flagging 3,454 vulnerabilities on clean inputs, maintains detection for only 4 functions under on-target attacks (CR = 0.12%), a 99.88% reduction in integrity.
+This pattern occurs under realistic access assumptions: surrogate-transfer attacks on black-box APIs (GPT-4o, GPT-5-mini) and cost-limited gray-box models (Qwen2.5-Coder-32B), supplemented by on-target optimization for white-box models (Llama3.1-8B, CodeAstra, StarCoder2-15B).
+Across this threat landscape, GPT-5-mini is a notable exception, exhibiting substantially lower attack success (44.04% union $\mathrm{ASR}_{\mathrm{cond}}$) compared to other models (87%–100%), suggesting broader invariance.
+Table 1 gives end-to-end post-attack metrics.
+Clean benchmark recall is not a security guarantee: an attacker needs only to evade the cases a detector correctly identifies, and our results show this is achievable via innocuous transformations (full carrier breakdown in Table 2).
+Table 1: Clean coverage and post-attack robustness (N=5000N{=}5000).
+Transfer: universal strings from Qwen2.5-Coder-14B surrogate applied without modification.
+On-target: GCG optimized directly on the target model.
+Models with only one row were evaluated under that regime only.
+Model Regime $\mathrm{TPR}_{\mathrm{clean}}$ (%) Union $\mathrm{ASR}_{\mathrm{cond}}$ (%) $\mathrm{CR}$ (%) $\Delta\mathrm{TPR}$ ($p$)
+Llama3.1-8B Transfer 31.04 99.68 0.32 30.94
+On-target 99.87 0.13 30.99
+CodeAstra Transfer 69.08 95.00 5.00 65.63
+On-target 99.88 0.12 69.01
+StarCoder2-15B Transfer 99.04 87.72 12.28 86.88
+On-target 90.39 9.61 89.52
+Qwen2.5-Coder-32B Transfer 22.42 99.38 0.62 22.28
+GPT-4o Transfer 46.00 87.00 13.00 40.02
+GPT-5-mini Transfer 73.62 44.04 55.96 32.42
+Figure 2: Non-GCG baselines.
+prompt injection tests output-format harness vulnerability (outside threat model); random_idsub tests code-understanding robustness (within threat model).
+Random Token Insertion on comment carriers achieves 13.2% vs.
+GCG’s 91.5%.
 Benign-class invariance.
-To verify that our edits induce targeted false negatives rather than indiscriminate label corruption, we evaluate attacks on a held-out set of 500 ground-truth benign functions.
+To verify that our edits induce targeted false negatives rather than label corruption, we evaluate attacks on a held-out set of 500 ground-truth benign functions.
 Across all models and attack methods, we observe 100% benign-class invariance: every function predicted as BENIGN under clean inference remains BENIGN after applying our transformations.
@@
 Model Identifier Sub. Comment ifdef macro dead-branch
-GPT-4o 31.09% 65.48% 71.61% 42.35% 19.09%
-GPT-5-mini 10.62% 18.34% 28.14% 21.57% 17.58%
 Qwen2.5-Coder-32B 26.94% 91.53% 98.84% 85.81% 63.15%
@@
 CodeAstra 4.72% 78.60% 42.59% 75.48% 72.12%
-We break down attack effectiveness by carrier under surrogate transfer (Table 2).
-Three patterns emerge.
-First, no single carrier dominates: ifdef achieves near-complete evasion on Llama3.1-8B (99.03%) and Qwen2.5-Coder-32B (98.84%) but achieves moderate success elsewhere.
-Second, preprocessor carriers often outperform identifier substitution, suggesting over-reliance on non-executable boundary tokens.
-Third, GPT-5-mini exhibits uniformly low ASR (10–28%), indicating broader invariance.
-Besides, Dead-branch carriers are also effective on several models, indicating limited dead code alignment.
+StarCoder2-15B 1.31% 4.00% 0.18% 86.83% 1.58%
+GPT-4o 31.09% 65.48% 71.61% 42.35% 19.09%
+GPT-5-mini 10.62% 18.34% 28.14% 21.57% 17.58%
+Table 2 breaks down attack effectiveness by carrier under surrogate transfer.
+No single carrier dominates across models: ifdef achieves near-complete evasion on Llama3.1-8B (99.03%) and Qwen2.5-Coder-32B (98.84%) but moderate success elsewhere.
+Preprocessor carriers consistently outperform identifier substitution, suggesting over-reliance on non-executable boundary tokens.
+GPT-5-mini exhibits uniformly low ASR (10–28%), indicating broader distributional invariance than other models.
+Dead-branch carriers are effective across multiple models, indicating limited dead-code resolution in current detectors.
 Model-specific patterns.
-Most models show high vulnerability to preprocessor carriers, with Llama3.1-8B and Qwen2.5-Coder-32B near-ceiling (85–99%).
-CodeAstra presents a distinct profile: identifier substitution transfers poorly (4.72%), while boundary carriers remain effective (42–79%).
-On-target optimization reverses this pattern (§5.4).
+Most models show high vulnerability to preprocessor carriers; Llama3.1-8B and Qwen2.5-Coder-32B are near-ceiling (85–99%).
+CodeAstra presents a distinct profile: identifier substitution transfers poorly (4.72%) while comment and preprocessor carriers remain effective (42–79%); on-target optimization reverses this (§5.4).
+Identifier substitution ASRs are computed over the full $\mathrm{TP}_{\mathrm{clean}}$ set, with compilation failures (9.1%; Table 4) counted as unsuccessful attacks (conservative estimate).
+StarCoder2-15B exhibits a single-carrier dependency: macro alone accounts for nearly all transfer evasion (86.83%), confirming that one vulnerable surface suffices to undermine overall robustness.
 Positional and functional sensitivity.
@@
 Function length shows weak correlation with evasion (Pearson $r\in[-0.17,-0.02]$), and CWE-level analysis reveals modest variation (Appendix D.1 and Figure 3).
-5.3 Sanitization Eliminates Carriers but Induces Prediction Drift
-Table 3: Sanitization eliminates targeted carriers (ASR $\rightarrow$ 0%) but induces prediction drift on clean inputs (see Table 10 for details).
+5.3 Defense Evaluation
+We evaluate three defenses of increasing strength, spanning inference-time preprocessing, inference-time ensembling, and training-time intervention.
 […]
-Qwen2.5-Coder-32B 91.53→0% 98.84→0% 85.81→0% +5.2%
-Llama3.1-8B 64.50→0% 99.03→0% 96.33→0% +8.6%
-CodeAstra 78.60→0% 42.59→0% 75.48→0% +0.1%
-Given the high success of comment- and preprocessor-based carriers, we evaluate a simple input-sanitization baseline that removes these carrier surfaces prior to inference.
-The pipeline (i) strips C/C++ line and block comments, and (ii) canonicalizes/removes preprocessor directives before passing code to the detector.
-We treat this as a carrier-elimination baseline rather than a semantics-aware defense.
-Table 3 shows that sanitization successfully drives ASR to zero for targeted carriers.
-However, it induces substantial prediction drift: stripping benign comments and preprocessor directives from unperturbed, ground-truth vulnerable functions increases clean detection rates by up to 8.6% (Llama3.1-8B: 1552→1685 true positives).
-This reveals that models utilize non-executable lexical context as decision-relevant features, indicating a lack of semantic invariance even absent adversarial perturbations (detailed drift analysis in Table 10, Appendix D).
-Moreover, sanitization is not a free lunch: comments and preprocessor directives appear in 57.0% and 12.8% of functions, respectively, meaning this defense rewrites a substantial fraction of inputs.
-We exclude dead-code elimination, as it risks mis-eliminating live code without compiler-level context (Tu et al., 2024).
-Furthermore, sanitization fails to address executable-looking carriers like identifier substitution and dead-branch insertions.
+Randomized smoothing.
+Adapting SmoothLLM Robey et al.
+(2024) to vulnerability detection, we generate K=5K{=}5 variants per input via random identifier renaming (up to 2 renames) and take majority vote.
+We evaluate this on the identifier substitution carrier under on-target GCG.
+ASR reductions are modest across all three evaluated models: Llama3.1-8B ($43.49\%\to 34.57\%$, 8.9-8.9 pp), CodeAstra ($85.41\%\to 82.71\%$, 2.7-2.7 pp), StarCoder2-15B ($27\%\to 14.63\%, 1.6-1.6$ pp).
+This confirms that GCG-optimized adversarial identifiers are robust to small random perturbations of the input space.
+Adversarial fine-tuning.
+To evaluate a training-time defense, we LoRA fine-tune Llama3.1-8B on carrier-diverse adversarial examples spanning all five carriers and re-optimize new universal strings (with a held-out seed) on the hardened model.
+Comment and preprocessor carriers drop sharply (comment: $42.01\%\to 0.16\%$; ifdef: $99.03\%\to 0.00\%$; macro: $65.46\%\to 16.91\%$), but identifier substitution ASR increases ($43.49\%\to 89.56\%$) and dead-branch remains high ($86.66\%\to 68.20\%$).
+Overall CR improves from 0.13% to 3.88%, while clean TPR drops from 31.04% to 24.72% (6.3-6.3 pp).
+Adversarial training thus closes specific carrier surfaces while potentially exposing others, and introduces a measurable robustness-accuracy trade-off.
+Cross-defense summary.
+No single defense is sufficient.
+Sanitization eliminates three carrier surfaces but leaves two intact and distorts the decision boundary.
+Smoothing yields marginal gains against optimized perturbations.
+Adversarial training provides the largest CR improvement but transfers unevenly across carriers.
+Effective mitigation likely requires combining complementary defenses across inference and training time; per-model breakdowns are in Appendix D.
 5.4 Transfer Attacks Succeed on Black-Box APIs; On-Target Reveals Complex Interactions
@@
 This models a low-barrier adversary: optimization is performed once, and resulting strings are reused across vulnerable functions and deployed detectors.
-Across black-box and operational black-box targets, transfer achieves 44.04%–99.03% union $\mathrm{ASR}_{\mathrm{cond}}$, establishing that API opacity alone does not prevent evasion under semantics-preserving perturbations.
-However, transferability is not uniform: GPT-5-mini exhibits materially lower success (44.04%) than other models (87–99%), indicating that robustness varies across model families even under identical carriers.
+Across black-box API and cost-limited gray-box targets, transfer achieves 44.04%–99.38% union $\mathrm{ASR}_{\mathrm{cond}}$, establishing that API opacity alone does not prevent evasion under syntax-preserving perturbations.
+However, transferability is not uniform: GPT-5-mini exhibits lower success (44.04%) than other models (87–99%).
+To test whether this reflects genuine robustness or insufficient surrogate capacity, we applied identifier substitution strings optimized on-target on Qwen2.5-Coder-32B (a stronger surrogate) to GPT-5-mini; ASR increases only marginally (10.62% $\to$ 13.4%), suggesting that the lower transferability is not an artifact of weak surrogate transfer.
 On-target optimization reveals divergent vulnerability profiles.
 While transfer attacks establish a practical lower bound on exploitability, they may underestimate the true vulnerability of a model.
-To quantify the gap between low-cost transfer and worst-case security risk, we perform on-target universal GCG for Llama3.1-8B and CodeAstra (accessible models).
-Table 4 compares transfer and on-target regimes across carriers.
-Table 4: Transfer vs. on-target attack success. Transfer uses universal strings from Qwen-14B surrogate; on-target optimizes directly on the target.
-Model Regime $\mathrm{ASR}_{\mathrm{cond}}$ (%)
-Idsub Comment ifdef macro Dead-br
-Llama3.1-8B Transfer 34.66 64.50 99.03 96.33 59.79
-On-target 43.49 42.01 99.03 65.46 86.66
-CodeAstra Transfer 4.72 78.60 42.59 75.48 72.12
-On-target 85.41 63.06 16.62 98.87 97.51
-For Llama3.1-8B, on-target optimization improves identifier substitution (34.66% $\rightarrow$ 43.49%) and dead-branch insertion (59.79% $\rightarrow$ 86.66%), but degrades comment (64.50% $\rightarrow$ 42.01%) and macro (96.33% $\rightarrow$ 65.46%).
-This suggests that universal strings optimized on a surrogate can outperform target-specific optimization for certain carriers, likely because surrogate-optimized universal strings exploit more transferable boundary-token sensitivities.
-CodeAstra exhibits a clear asymmetry. Transfer achieves 4.72% identifier success but 42.59% for ifdef. On-target optimization inverts this: identifier reaches (85.41%) while ifdef collapses (16.62%).
+To quantify the gap between low-cost transfer and worst-case security risk, we perform on-target universal GCG forLlama3.1-8B, CodeAstra, and StarCoder2-15B (white-box models); full per-carrier breakdown in Table 7, Appendix C.2).
+StarCoder2-15B shows consistent gains under on-target optimization across most carriers (e.g., comment: 4.00% $\to$ 69.59%), reducing CR from 12.28% to 9.61%.
+For Llama3.1-8B, on-target optimization improves identifier substitution ($34.66\%\to 43.49\%$) and dead-branch ($59.79\%\to 86.66\%$) but degrades comment ($64.50\%\to 42.01\%$) and macro ($96.33\%\to 65.46\%$).
+Universal strings optimized on a surrogate can outperform target-specific optimization for certain carriers, likely because cross-model training discovers more transferable boundary-token sensitivities.
+CodeAstra exhibits a clear asymmetry.
+Transfer achieves 4.72% identifier success but 42.59% for ifdef.
+On-target optimization inverts this: identifier reaches 85.41% while ifdef collapses (16.62%).
 We hypothesize two mechanisms.
 For identifier substitution, CodeAstra’s loss landscape provides smooth gradients: GCG efficiently discovers variable-name perturbations that flip predictions.
-For ifdef, on-target GCG likely encounters optimization difficulty, so the optimizer converges to poor local minima that fail to exploit the vulnerabilities that surrogate-optimized strings (accidentally) discovered through multi-model training.
-This suggests that gradient-based optimization success depends not only on model vulnerabilities but also on the optimizability of the attack surface: some carrier spaces may be inherently harder to optimize even when vulnerabilities exist.
+For #ifdef, on-target GCG likely encounters optimization difficulty, converging to poor local minima that fail to exploit the vulnerabilities that surrogate-optimized strings discovered through multi-model training.
+This suggests gradient-based optimization success depends not only on model vulnerabilities but also on the optimizability of the attack surface.
+A related observation on Qwen2.5-Coder-14B (identifier substitution converges smoothly while comment insertion oscillates; Appendix C.3) provides indirect support; direct loss-landscape analysis on CodeAstra #ifdef is left to future work.
 Qwen2.5-Coder-32B: Surrogate capacity and on-target upper bound.
-For Qwen2.5-Coder-32B (operational black-box), we ablate identifier substitution across surrogate capacities and on-target optimization (Table 5).
+For Qwen2.5-Coder-32B (cost-limited gray-box), we ablate identifier substitution across surrogate capacities and on-target optimization (Table 8, Appendix C.2).
 Transferability is non-zero even for small surrogates (0.5B/1.5B:  6-7%), but increases with capacity (14B: 26.94%), indicating a representation-capacity threshold.
-On-target universal GCG achieves 92.50%, providing a clear upper bound when gradients are available.
-This gap (26.94% $\rightarrow$ 92.50%) quantifies the security improvement from restricting adversarial access: surrogate transfer yields a conservative lower bound under practical compute constraints, while on-target optimization demonstrates worst-case exploitability.
-Table 5: Identifier substitution on Qwen2.5-Coder-32B: effect of surrogate capacity and on-target optimization. All settings use the same injection site, identifier constraints, and $L_{\max}$.
-Setting $\mathrm{ASR}_{\mathrm{cond}}$ (%) Flipped / $|\mathbf{TP}_{\mathrm{clean}}|$
-Surrogate transfer (Qwen2.5-Coder-0.5B $\rightarrow$ 32B) 6.87 77 / 1121
-Surrogate transfer (Qwen2.5-Coder-1.5B $\rightarrow$ 32B) 6.42 72 / 1121
-Surrogate transfer (Qwen2.5-Coder-14B $\rightarrow$ 32B) 26.94 302 / 1121
-On-target universal GCG (Qwen2.5-Coder-32B) 92.50 1038 / 1121
-Implications.
-These findings highlight three key insights.
-First, transferability amplifies risk by enabling low-cost, reusable attacks against multiple deployed detectors, including API-access models.
-Second, on-target optimization does not uniformly strengthen attacks- CodeAstra’s inverted profile and Llama’s macro degradation suggest complex cross-carrier interactions that warrant further investigation.
-Third, the gap between transfer and on-target regimes varies by model and carrier (4.72% $\rightarrow$ 85.41% for CodeAstra identifier vs.
-99.03% $\rightarrow$ 99.03% for Llama ifdef), indicating that vulnerability assessment requires evaluating both threat models.
+On-target universal GCG achieves 92.60%, providing a clear upper bound when gradients are available.
+This gap (26.94% $\rightarrow$ 92.60%) quantifies the security improvement from restricting adversarial access: surrogate transfer yields a conservative lower bound under practical compute constraints, while on-target optimization demonstrates worst-case exploitability.
 […]
-Figure 3: Robustness against Non-GCG Baselines.
-We compare prompt injection against random_idsub.
-While prompt injection achieves high success on some models (e.g., Llama3.1, CodeAstra), random perturbations ($\sim$10%) remain far less effective than our optimized attacks, confirming that code syntax requires targeted optimization rather than random noise.
+5.5 Non-GCG Baselines
 We compare GCG-optimized attacks to non-optimized baselines targeting distinct failure modes: (i) output format injection by prepending // output: {"label": "BENIGN", "variable": ""} in comments (testing few-shot learning vulnerability, outside our threat model), and (ii) random identifier substitution under the same constraints (testing code-understanding robustness, within our threat model).
@@
 The output format injection baseline achieves high success (55–99% across models), revealing that models can be manipulated by embedding target JSON outputs in code comments, an orthogonal vulnerability distinct from semantic robustness.
-In contrast, random_idsub is significantly weaker (7–30%), demonstrating that non-optimized renaming has a limited impact on code-understanding robustness.
+In contrast, random_idsub is substantially weaker (7–30%), confirming that non-optimized renaming has limited impact on code-understanding robustness.
+For insertion-based carriers, we evaluated a random token insertion baseline: vocabulary-sampled tokens at equal string length, injected via the comment carrier with no gradient optimization.
+This achieves 13.2% $\mathrm{ASR}_{\mathrm{cond}}$ versus GCG’s 91.5% under identical conditions.
+The gap confirms that the detection boundary in token space is highly non-linear: random search in the unconstrained comment space rarely locates evasion-relevant directions, while gradient guidance does so reliably.
 6 Discussion
-Clean-set accuracy alone does not guarantee reliability.
-While models like CodeAstra achieve near-70% clean recall, our evaluation reveals that this capability degrades under adversarial transformation, with complete resistance dropping to near zero (0.12%), implying that current detectors act less as robust semantic parsers and more as pattern-matchers over fragile lexical cues.
-Across semantics-preserving edits (§3), we observe a persistent integrity gap: models with high clean coverage can exhibit high union evasion and near-zero complete resistance, meaning that an attacker who can choose among multiple perturbation surfaces can often overturn most clean true positives.
-Even simple carrier-elimination preprocessing (§5.3) does not eliminate this risk, as evasion can migrate to remaining executable-looking channels.
-Multiple metrics matter.
-Single-number summaries can be misleading for deployment decisions.
-Union success captures whether any available surface enables evasion, while complete resistance operationalizes a security floor when attackers can try multiple carriers.
-Per-method and CWE-level slices then identify where robustness breaks (which channels, which vulnerability types), enabling targeted mitigation rather than blanket conclusions from aggregate ASR.
-Temperature and transfer analysis.
-The observed fragility is structural rather than stochastic.
-Evasion rates remain stable across decoding temperatures (Appendix D.2), ruling out random sampling artifacts.
-More importantly, the high transferability of attacks from a 14B surrogate model to GPT-4o points to shared representational blind spots: distinct models appear to over-index on identical non-executable tokens (e.g., #ifdef guards) while ignoring the invariant executable logic.
-Decoupling of Domain Capability and Robustness.
-The performance profile of CodeAstra reveals a meaningful insight: although task-specific fine-tuning boosts baseline detection, it does not guarantee robustness.
-CodeAstra exhibits high vulnerability in specific channels (e.g., macro insertion) despite high clean accuracy.
-This suggests that domain specialization can coexist with-or mask-structural fragility, emphasizing that fine-tuned models require integrity evaluations just as rigorous as general-purpose ones.
-Implications for Deployment and Defense.
-Our findings suggest three concrete directions for hardening detectors.
-First, training on carrier-diverse adversarial examples (spanning identifier, comment, and preprocessor surfaces) may enforce broader invariance than single-method defenses.
-Second, Complete Resistance should join clean accuracy as a co-equal deployment metric- a detector is only as robust as its weakest carrier.
-Third, ensemble approaches combining sanitization (for boundary carriers) with robust base models may offer defense-in-depth where neither alone suffices.
-Limitations and Future Directions.
-We focus on C/C++ functions with a fixed prompting protocol; extending to other languages and detection formulations would strengthen generality.
-While Tier 2 validation (differential AST analysis) scales to 5,000 functions and Tier 3 (compilation-based) provides strong proof on 1,434 functions (Appendix A.3), absolute behavioral equivalence requires full execution environments unavailable in function-level datasets.
-Future work should explore adversarial fine-tuning on semantics-preserving carriers, though computational cost and overfitting risks motivate investigating lightweight defense wrappers (Madry et al., 2019; Varshney et al., 2024).
+Clean recall as a misleading proxy.
+Under the five syntax-preserving carriers evaluated, even high-accuracy detectors exhibit near-zero Complete Resistance, indicating reliance on non-executable tokens rather than program semantics.
+Union ASR captures whether any carrier enables evasion, while CR operationalizes a security floor across carriers.
+We note that several models (CodeAstra: 70% FPR, GPT-5-mini: 54% FPR on benign inputs; Appendix B.3) are questionable as standalone deployment gates even before adversarial pressure.
+Our CR metric complements broader multi-objective evaluation frameworks such as SecLens-R (Halder et al., 2026), adding a robustness dimension to stakeholder-aware vulnerability detection assessment.
+Defense landscape.
+No evaluated defense is sufficient in isolation (§5.3).
+Input sanitization eliminates comment and preprocessor carriers but induces prediction drift of up to 8.6% and leaves identifier substitution and dead-branch carriers intact.
+Randomized smoothing yields only marginal reductions (1.6–8.9 pp).
+Adversarial fine-tuning provides the largest CR improvement (0.13% $\to$ 3.88%) but shifts vulnerability to the identifier carrier and reduces clean TPR by 6.3 pp.
+These results suggest that effective mitigation requires combining complementary defenses across inference and training time; carrier-diverse adversarial training (Madry et al., 2019) with broader carrier coverage and purification wrappers (Mu et al., 2025) are concrete next steps.
+Limitations.
+We focus on C/C++ with a fixed prompting protocol; extension to other languages would strengthen generality.
+Our three-tier validation certifies syntax- and compilation-level preservation only; formal runtime equivalence requires execution environments unavailable in function-level datasets (Appendix A.3).
+The optimized universal strings (particularly for identifier substitution) are often unnatural and may be flagged by style linters or repository naming policies; the attack success rates reported here assume no such pre-screening.
+Multi-seed analysis covers on-target identifier substitution across three models; broader carrier and regime coverage remains future work.
 7 Conclusion
-On a unified C/C++ benchmark (N=5000N{=}5000), we present an integrity evaluation of LLM-based vulnerability detectors under semantics-preserving code perturbations, evaluating five carriers spanning four insertion families (identifier substitution, comments, preprocessor insertions, and dead-branch insertion) and quantifying robustness via union success and complete resistance.
-Across models, we observe widespread failures of semantic invariance: many vulnerabilities that are correctly detected on clean inputs can be flipped to false negatives via behavior-equivalent edits, and a single universal, carrier-constrained string optimized on a surrogate can transfer unchanged to black-box API targets.
-Sanitization can remove some insertion surfaces, but it rewrites many inputs and leaves executable-looking channels vulnerable.
-These findings collectively suggest that clean benchmark performance alone is insufficient as a security guarantee, and that integrity-oriented evaluations and defenses should explicitly target robustness to behavior-preserving code changes without resorting to overly conservative task-altering preprocessing.
-Impact Statement
-This work evaluates the robustness of LLM-based vulnerability detectors under semantics-preserving adversarial perturbations.
-We discuss both the positive contributions and potential risks of this research.
-Positive Impacts.
-Our findings reveal essential blind spots in current LLM-based code security tools before they are exploited in the wild.
-By systematically characterizing failure modes and introducing metrics such as Complete Resistance, we provide actionable diagnostics for practitioners to harden detection pipelines.
-We believe responsible disclosure of these weaknesses, consistent with standard practices in security research, enables the community to develop more robust defenses.
-Potential Risks and Mitigations.
-As with most security research, the techniques described could theoretically be misused to evade vulnerability detection.
-However, we note that: (1) the attack surfaces we study (identifier renaming, comment insertion, preprocessor directives) are already well-known code transformations; (2) we do not release optimized adversarial strings or attack toolkits; and (3) the primary barrier to exploiting vulnerabilities lies in discovering them, not in evading detection.
-We believe the defensive value of exposing these weaknesses substantially outweighs the marginal risk of informing attackers about evasion strategies they could likely discover independently.
+On a unified C/C++ benchmark (N=5000N{=}5000), we evaluate six LLM-based vulnerability detectors under five syntax-preserving (Tree-sitter-validated) and largely compilation-preserving (gcc-validated) carrier variants.
+Complete Resistance falls to at most 13% for five of six models, and a universal string optimized on a 14B surrogate transfers unchanged to black-box APIs including GPT-4o (87% union $\mathrm{ASR}_{\mathrm{cond}}$).
+Among defenses, sanitization addresses only comment and preprocessor surfaces, and randomized smoothing yields marginal reductions (1.6–8.9 pp).
+Adversarial fine-tuning improves CR from 0.13% to 3.88% but introduces a robustness-accuracy trade-off.
+Multi-seed analysis across three models confirms that the observed exploitability is systematic rather than seed-dependent.
+These results indicate that clean benchmark recall is not a proxy for deployment robustness; integrity-oriented evaluations and training-time defenses must target robustness to syntax-preserving changes.
 Appendix A Detailed Threat Model
 A.1 Target Models Details
-We evaluate the proposed attacks across five state-of-the-art LLMs, categorized by their access regimes:
+We evaluate the proposed attacks across six target LLMs and one surrogate, categorized by their access regimes:
 For reproducibility, we explicitly set temperature=0.0 for GPT-4o to ensure deterministic decoding, while GPT-5-mini was evaluated under its mandatory default settings, as the API does not currently support sampling parameter adjustment.
@@
 White-box Llama 3.1 meta-llama/Llama-3.1-8B-Instruct
-White-box CodeAstra rootxhacker/CodeAstra-7B (Ganesan, 2024)
+White-box CodeAstra rootxhacker/CodeAstra-7B Ganesan (2024)
+White-box StarCoder2 bigcode/starcoder2-15b-instruct-v0.1
 White-box Qwen2.5-Coder Qwen/Qwen2.5-Coder-14B-Instruct (Surrogate)
@@
 Black-box (API) GPT-5-mini gpt-5-mini-2025-08-07
-Operational Black-box Qwen2.5-Coder Qwen/Qwen2.5-Coder-32B-Instruct
+Cost-limited Gray-box Qwen2.5-Coder Qwen/Qwen2.5-Coder-32B-Instruct
 A.2 Vocabulary Constraints
@@
   • Identifiers and Dead-branch. Restricted to [A-Za-z_][A-Za-z0-9_]*.
+  • Comment carrier. We restrict to printable ASCII excluding sequences that could terminate the comment block (*/, newlines).
   • Macro Bodies and Guards. We use a filtered printable ASCII set $\mathcal{A}_{\text{macro}}$ that excludes characters capable of terminating string literals or lines prematurely:
@@
 […]
-A.3 Semantics Preservation Validation
-We validate semantics preservation through a two-tiered approach: (1) Constraint-based Validation to verify method-specific injection rules, and (2) Differential Syntactic Analysis using Tree-sitter to ensure that perturbations do not introduce new syntax errors compared to the original code.
-This approach is robust to the incomplete nature of function-level datasets (e.g., missing headers) while strictly flagging syntax violations.
+A.3 Syntax Preservation and Compilation Validation
+We validate syntax and compilation preservation through a three-tiered approach: (1) Constraint-based Validation to verify method-specific injection rules, (2) Differential Syntactic Analysis using Tree-sitter to ensure perturbations do not introduce new syntax errors compared to the original code (applied on all 5,000 functions), and (3) Compilation-based Validation via gcc on a 1,434-function compilable subset.
+All three tiers certify syntactic and compilation-level preservation only; formal runtime-equivalence would require full execution environments unavailable in function-level datasets.
 Tier 1: Constraint-based Validation. For each attack method, we verify rigorous structural constraints:
@@
 This cross-model consistency confirms that our adversarial perturbations- whether applied to small or large test sets-preserve the syntactic structure of the original code, effectively ruling out malformed injections as a cause for evasion.
-Table 6: Semantics Preservation Rates across Models.
+Table 3: Syntax Preservation Rates across Models (Tree-sitter Tier).
 We report two metrics: (1) Constraint Sat.
@@
 Fraction of samples where the attack introduces no new AST error nodes ($E_{adv}\leq E_{orig}$).
+These rates certify syntactic preservation only; compilation-level validation is reported separately in Table 4.
 High validity across diverse architectures confirms the stability of our attack generation.
@@
   3. 3.
-    Evaluate on CodeAstra (highest TPR=69.08%) → 1,434 correctly detected (71.7%)
+    Evaluate on CodeAstra (TPR=69.08%; primary white-box target) $\rightarrow$ 1,434 correctly detected (71.7%)
   4. 4.
@@
 Table 4 shows compilation preservation rates.
-Additive carriers achieve perfect 100% preservation (1,434/1,434), providing compiler-verified proof of semantics-preservation by construction.
+Comment, preprocessor, and dead-branch carriers achieve perfect 100% compilation-preservation (1,434/1,434), confirming their non-interference at the compilation level; this is consistent with their construction (comments and preprocessor directives are stripped before compilation; dead-branch is never executed).
 Identifier Substitution achieves 90.9% (1,303/1,434).
-The 131 failures stem from collisions with injected header symbols (uint8_t, size_t, printf, malloc, etc.) and are excluded from ASR calculations.
+The 131 failures stem from collisions with injected header symbols (uint8_t, size_t, printf, malloc, etc.).
+These cases are counted as unsuccessful attacks in ASR computation (included in the $\mathrm{TP}_{\mathrm{clean}}$ denominator but not as flipped), providing a conservative ASR estimate.
 Table 4: Compilation Preservation under Transfer Attacks (N=1,434N{=}1,434). CodeAstra true positives attacked with universal strings from Qwen-14B (Table 2). Validation via gcc -fsyntax-only.
@@
 Data sources and target language.
-Unified-VUL-N contains vulnerable functions written in C/C++ and isconstructed by pooling the vulnerable subsets from PrimeVul, BigVul, and DiverseVul.The goal is to evaluate robustness under semantics-preserving transformations on asingle, fixed benchmark with controlled composition.
+Unified-VUL-N contains vulnerable functions written in C/C++ and is constructed by pooling the vulnerable subsets from PrimeVul, BigVul, and DiverseVul.
+The goal is to evaluate robustness under syntax-preserving transformations on a single, fixed benchmark with controlled composition.
 Normalization for cross-source deduplication.
@@
 To guarantee sufficient headroom for injected strings and to standardize the context budget across models, we filter functions by code-only token length using the Qwen2.5-Coder-14B tokenizer.
-Specifically, we require the code-only length to be at most 4096 tokens, where “code-only” refers to the function text used as input code (excluding any additional prompt or JSON schema wrappers).
+Specifically, we require the code-only length to be at most 4096 tokens, where "code-only" refers to the function text used as input code (excluding any additional prompt or JSON schema wrappers).
 This constraint ensures that our injected variants remain within model context limits under the shared evaluation prompt.
@@
 Sanitized FPR is computed after applying the same sanitization pipeline as in §5.3 (comment stripping and preprocessor removal); we report $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ as the change in false-positive rate (percentage points) from clean to sanitized inputs.
-Benign-class Invariance Rate is the fraction of benign functions that are predicted as BENIGN under clean inference and remain BENIGN after applying our semantics-preserving attacks (union over all attack methods).
+Benign-class Invariance Rate is the fraction of benign functions that are predicted as BENIGN under clean inference and remain BENIGN after applying our syntax-preserving attacks (union over all attack methods).
 This metric directly measures whether our attacks preserve benign-class predictions, which is aligned with our attack objective (evading vulnerability detection without introducing false positives).
@@
 Table 5 reports per-model clean FPR, sanitized FPR (and $\Delta\mathrm{FPR}_{\mathrm{sanit}}$), and benign-class invariance rate.
-All models achieve 100% benign-class invariance rate, demonstrating that our semantics-preserving attacks preserve benign-class predictions: all functions that are correctly classified as BENIGN under clean inference remain BENIGN after attack.
+All models achieve 100% benign-class invariance rate, demonstrating that our syntax-preserving attacks preserve benign-class predictions: all functions that are correctly classified as BENIGN under clean inference remain BENIGN after attack.
 Table 5: Benign-side effects and invariance.
@@
 $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ is the change (percentage points) after sanitization.
-Benign-class Invariance is computed on benign functions predicted as BENIGN under clean inference and remains BENIGN after applying the union of semantics-preserving attacks.
+Benign-class Invariance is computed on benign functions predicted as BENIGN under clean inference and remains BENIGN after applying the union of syntax-preserving attacks.
 Model $N_{\mathrm{benign}}$ Clean FPR (%) Sanitized FPR (%) $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ ($p$) Benign Invariance (%)
@@
 To ensure fair cross-model comparisons, we standardized length filtering using the Qwen2.5-Coder-14B tokenizer to maintain an identical evaluation set across all targets.
-Per-model filtering would result in inconsistent subsets, making metrics like $TPR_{clean}$ and $ASR_{cond}$ non-comparable.
+Per-model filtering would result in inconsistent subsets, making metrics like $\mathrm{TPR}_{\mathrm{clean}}$ and $\mathrm{ASR}_{\mathrm{cond}}$ non-comparable.
 Correlation analysis on 500 random samples confirms that the Qwen tokenizer is highly representative: it achieves a Pearson correlation of $r=0.9998$ with Llama-3.1-8B and $r=0.9941$ with Mistral-7B-Instruct-v0.2, which is the original model of CodeAstra.
@@
 Cost metrics.
-We track (i) forward passes for gradients ($m_{c}$ per step during curriculum, $m_{\mathrm{train}}$ per step after balance), (ii) forward passes for candidate evaluation ($m_{c}\cdot\lceil\texttt{search\_width}/\texttt{batch\_size}\rceil=m_{c}\cdot 8$ per step with our default settings), and (iii) wall-clock time (total optimization time including GPU memory transfers and batched operations).
+We track (i) forward passes for gradients ($m_{c}$ per step during curriculum, $m_{\mathrm{train}}$ per step after balance), (ii) forward passes for candidate evaluation ($m_{c}\cdot\lceil\texttt{search\_width}/\texttt{batch\_size}\rceil=m_{c}\cdot 32$ per step with our default settings), and (iii) wall-clock time (total optimization time including GPU memory transfers and batched operations).
 Optional per-step model.generate() calls are counted separately and can be disabled via enable_step_generate=False to reduce latency.
@@
 Positional Sensitivity.
-Table 9 compares the $ASR_{cond}$ for different comment-carrier insertion sites.
+Table 6 compares the $\mathrm{ASR}_{\mathrm{cond}}$ for different comment-carrier insertion sites.
 We observe that Trailing comments (placed after the function body) often exhibit high effectiveness, notably for Llama3.1-8B and CodeAstra, where success rates significantly exceed those of the original Leading header comments.
 However, Middle placements (interleaved with the logic) often show lower effectiveness across all models.
-This suggests that the model’s contextual focus on interleaved logic may partially mitigate the influence of semantics-preserving perturbations compared to injections at the function boundaries.
+This suggests that the model’s contextual focus on interleaved logic may partially mitigate the influence of syntax-preserving perturbations compared to injections at the function boundaries.
 Functional Role Sensitivity.
@@
 Table 6: Ablation results for different injection sites and functional roles.
-$ASR_{cond}$ is reported as the fraction of originally correctly detected vulnerable functions ($TP_{clean}$) that were flipped to BENIGN.
-Model Family Variant Flipped / $TP_{clean}$ $ASR_{cond}$ (%)
+$\mathrm{ASR}_{\mathrm{cond}}$ is reported as the fraction of originally correctly detected vulnerable functions ($\mathrm{TP}_{\mathrm{clean}}$) that were flipped to BENIGN; Qwen2.5-Coder-32B uses transfer strings; Llama3.1-8B and CodeAstra use on-target strings.
+Model Family Variant Flipped / $\mathrm{TP}_{\mathrm{clean}}$ $\mathrm{ASR}_{\mathrm{cond}}$ (%)
 Qwen2.5-Coder-32B Comment Head (Original) 1026 / 1121 91.53
@@
 Body 1063 / 3454 30.78
-Appendix D Extended Sanitization Analysis
+C.2 Transfer vs. On-Target Breakdown
+Table 7 provides the full carrier-level comparison of transfer and on-target attack success across three white-box models.
+Table 8 shows surrogate-capacity ablation for Qwen2.5-Coder-32B identifier substitution.
+Table 7: Transfer vs. on-target attack success. Transfer uses universal strings from Qwen-14B surrogate; on-target optimizes directly on the target.
+Model Regime $\mathrm{ASR}_{\mathrm{cond}}$ (%)
+Idsub Comment ifdef macro Dead-br
+Llama3.1-8B Transfer 34.66 64.50 99.03 96.33 59.79
+On-target 43.49 42.01 99.03 65.46 86.66
+CodeAstra Transfer 4.72 78.60 42.59 75.48 72.12
+On-target 85.41 63.06 16.62 98.87 97.51
+StarCoder2-15B Transfer 1.31 4.00 0.18 86.83 1.58
+On-target 16.27 69.59 5.45 85.25 32.86
+Table 8: Identifier substitution on Qwen2.5-Coder-32B: effect of surrogate capacity and on-target optimization. All settings use the same injection site, identifier constraints, and $L_{\max}$.
+Setting $\mathrm{ASR}_{\mathrm{cond}}$ (%) Flipped / $|\mathbf{TP}_{\mathrm{clean}}|$
+Surrogate transfer (Qwen2.5-Coder-0.5B $\rightarrow$ 32B) 6.87 77 / 1121
+Surrogate transfer (Qwen2.5-Coder-1.5B $\rightarrow$ 32B) 6.42 72 / 1121
+Surrogate transfer (Qwen2.5-Coder-14B $\rightarrow$ 32B) 26.94 302 / 1121
+On-target universal GCG (Qwen2.5-Coder-32B) 92.60 1038 / 1121
+C.3 GCG Loss Trajectories
+To support the hypothesis in §5.4 about carrier-level optimization difficulty, we ran GCG universal optimization on the Qwen2.5-Coder-14B surrogate for m=10m{=}10 vulnerable samples over 500 steps, recording per-step loss for both Identifier Substitution and Comment Insertion.
+After a curriculum phase completing at step 10, Identifier Substitution loss peaks at 2.89, then descends monotonically over ${\sim}300$ steps ($2.75\to 2.39\to 1.98\to 1.64$), stabilizing at 1.251.251.301.30.
+Comment Insertion, under the same protocol, spikes to 7.12 once all 10 samples are loaded, briefly drops to ${\sim}4.62$, then oscillates between $4.62$ and $4.75$ for the remaining 400 steps without further progress.
+These trajectories are consistent with Table 7: the constrained token space of valid C identifiers acts as a structural regularizer, aligning cross-sample gradients and enabling smooth convergence, whereas the unconstrained comment space allows gradients from diverse samples to interfere destructively, trapping the optimizer in a high-loss plateau.
+Appendix D Extended Defense Analysis
+Table 9: Sanitization eliminates targeted carriers (ASR $\rightarrow$ 0%) but induces prediction drift on clean inputs (see Table 10 for details).
+[…]
+[…]
+[…]
+[…]
 Table 10 details prediction drift when applying sanitization (stripping comments and preprocessor directives) to clean, unperturbed vulnerable functions.
@@
 As shown in Table 10, the detection boundary is highly unstable.
-For Llama3.1-8B, simply removing developer comments (which have no effect on program execution) causes the model to ”discover” 133 previously missed vulnerabilities—an 8.6% increase.
+For Llama3.1-8B, simply removing developer comments (which have no effect on program execution) causes the model to "discover" 133 previously missed vulnerabilities—an 8.6% increase.
 This suggests that the model’s clean accuracy is partly contingent on specific lexical distributions rather than robust semantic understanding.
@@
 ifdef 71.61% 74.17% $\pm$ 0.64%
-Appendix E Prompt templates
+D.3 Randomized Smoothing Details
+We adapt SmoothLLM Robey et al.
+(2024) by generating K=5K{=}5 perturbed copies of each input via random identifier renaming (uniformly sampling up to 2 identifiers per copy and replacing each with a random valid C identifier).
+The final prediction is the majority vote across the $K$ outputs.
+We evaluate on the identifier substitution carrier under on-target GCG strings, as this carrier is the most directly affected by identifier-level perturbations.
+Table 13: Randomized smoothing defense results on identifier substitution (on-target GCG). K=5K{=}5 variants, up to 2 renames.
+Model Original ASR Defended ASR $\Delta$ ($p$)
+Llama3.1-8B 43.49% 34.57% 8.9-8.9
+CodeAstra 85.41% 82.71% 2.7-2.7
+StarCoder2-15B 16.27% 14.63% 1.6-1.6
+The modest reductions indicate that GCG-optimized identifier strings occupy robust regions of the token space: small random perturbations do not displace the adversarial signal sufficiently to restore correct detection.
+D.4 Adversarial Fine-Tuning Details
+[…]
+After fine-tuning, we re-optimize new universal GCG strings on the hardened model using a held-out seed (seed=99) to avoid evaluating on memorized attack strings.
+Table 14: Adversarial fine-tuning results on Llama3.1-8B. New GCG strings re-optimized on the hardened model (seed=99).
+Carrier Original ASR Adv-Trained ASR $\Delta$ ($p$)
+Identifier Sub. 43.49% 89.56% +46.1+46.1
+Comment 42.01% 0.16% 41.9-41.9
+ifdef 99.03% 0.00% 99.0-99.0
+macro 65.46% 16.91% 48.6-48.6
+Dead-branch 86.66% 68.20% 18.5-18.5
+Union ASR 99.87% 96.12% 3.8-3.8
+CR 0.13% 3.88% +3.8+3.8
+Clean TPR 31.04% 24.72% 6.3-6.3
+Adversarial training effectively suppresses the three carrier families present in the training data as non-executable insertions (comment, ifdef, macro).
+However, identifier substitution ASR increases substantially, suggesting that the fine-tuning process shifts the decision boundary in a way that makes the model more sensitive to variable-name perturbations.
+The clean TPR reduction (6.3-6.3 pp) reflects an over-correction toward the VULNERABLE class on adversarial-looking inputs, consistent with known robustness-accuracy trade-offs in adversarial training Madry et al.
+(2019).
+Appendix E Prompt Transferability
+Our threat model (§3.1) does not assume adversarial control of the inference prompt; the adversary modifies only the code input.
+To verify that adversarial strings generalize beyond the specific prompt template used during optimization (Appendix F), we evaluated our best-performing carrier attacks against an entirely unseen zero-shot prompt:
+Is the following code vulnerable?
+{func}
+No re-optimization was performed; we applied frozen $\sigma$ values optimized under the original structured JSON prompt.
+Results.
+#ifdef insertion achieves 99.73% $\mathrm{ASR}_{\mathrm{cond}}$ and comment insertion achieves 97.23% $\mathrm{ASR}_{\mathrm{cond}}$ under this unseen prompt.
+These results confirm that the adversarial strings exploit the model’s internal feature representations rather than overfitting to a specific prompt template, consistent with the cross-model transfer results in §5.4.
+Prompt-diverse robustness testing of this kind is identified as a design-scope requirement for future evaluations.
+Appendix F Prompt templates
 We used the following prompt template (implemented in code as ’get_prompt_template()’):
@@
 JSON:
-Appendix F Attack Carrier Templates
+Appendix G Attack Carrier Templates
 To ensure full reproducibility, we provide the exact code templates used for each carrier family, corresponding to the implementation logic in our evaluation pipeline.
 In the templates below, [TRIGGER] denotes the GCG-optimized adversarial string (referred to as {optim_str} in the codebase), and [ORIGINAL_CODE] denotes the victim function text.
-F.1 Comment Carrier
+G.1 Comment Carrier
 We inject a Doxygen-style multi-line comment block immediately preceding the function definition.
-The adversarial string is embedded within the @commit_hash metadata field, acting as a ”distractor” in the non-executable header.
+The adversarial string is embedded within the @commit_hash metadata field, acting as a "distractor" in the non-executable header.
  * @brief Function implementation
@@
 [ORIGINAL_CODE]
-F.2 Macro Insertion
+G.2 Macro Insertion
 We prepend a macro definition before the function.
@@
 [ORIGINAL_CODE]
-F.3 Dead-Branch Insertion
+G.3 Dead-Branch Insertion
 Unlike the preprocessor carriers, this payload is injected inside the function body.
@@
         "benign_[TRIGGER]";
-F.4 #ifdef Insertion
+G.4 #ifdef Insertion
 We wrap a benign log statement containing the adversarial string within an #ifdef DEBUG_MODE block.
@@
 [ORIGINAL_CODE]
-F.5 Identifier Substitution
+G.5 Identifier Substitution
 For the identifier substitution attack (not template-based but replacement-based), we identify the target variable $v$ (via the model’s "variable" attribution or frequency analysis) and apply a global string replacement: